### v1

# Beyond Standard MoE: Mixture of Latent Experts for Resource-Efficient Language Models

Abstract
Mixture of Experts (MoE) has emerged as a pivotal architectural paradigm for efficient scaling of Large Language Models (LLMs), operating through selective activation of parameter subsets for each input token. Nevertheless, conventional MoE architectures encounter substantial challenges, including excessive memory utilization and communication overhead during training and inference, primarily attributable to the proliferation of expert modules. In this paper, we introduce Mixture of Latent Experts (MoLE), a novel parameterization methodology that facilitates the mapping of specific experts into a shared latent space. Specifically, all expert operations are systematically decomposed into two principal components: a shared projection into a lower-dimensional latent space, followed by expert-specific transformations with significantly reduced parametric complexity. This factorized approach substantially diminishes parameter count and computational requirements. Beyond the pretraining implementation of the MoLE architecture, we also establish a rigorous mathematical framework for transforming pre-trained MoE models into the MoLE architecture, characterizing the sufficient conditions for optimal factorization and developing a systematic two-phase algorithm for this conversion process. Our comprehensive theoretical analysis demonstrates that MoLE significantly enhances computational efficiency across multiple dimensions while preserving model representational capacity. Empirical evaluations corroborate our theoretical findings, confirming that MoLE achieves performance comparable to standard MoE implementations while substantially reducing resource requirements.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse natural language processing tasks (Bommasani et al., 2021; Zhuang et al., 2020), from text generation (Achiam et al., 2023; Dubey et al., 2024) to complex reasoning (Guo et al., 2025). As these models scale to increasingly larger parameter spaces, the Mixture of Experts (MoE) architecture (Jacobs et al., 1991; Jordan and Jacobs, 1994) has emerged as a promising paradigm for efficiently scaling model capacity without proportionally increasing computational costs. By selectively activating only a subset of parameters for each input token, MoE models achieve parameter efficiency while maintaining manageable inference latency.
Despite their theoretical and empirical advantages, standard MoE architectures (Dai et al., 2024) face significant practical limitations that inhibit broader deployment. These models require substantial memory resources to store parameters across numerous expert modules in Feed-Forward Network (FFN) layers and create communication bottlenecks during distributed training due to all-to-all data transfers. These challenges become increasingly pronounced as models scale to hundreds of experts, potentially limiting their applicability in resource-constrained environments.
Through systematic investigation of parameter utilization in MoE architectures, we discover substantial redundancy within the FFN layers of current MoE models. By analyzing Qwen1.5-MoE-A2.7B (Team, 2024), we reveal that a significant proportion of parameters in its FFN layers can be effectively approximated through lower-dimensional representations without compromising model performance. This empirical observation motivates a fundamental rethinking of expert parameterization in neural architectures.
Figure 1: Architectural comparison between MoE and MoLE in the feed-forward network (FFN) layer. In both diagrams, $N$ denotes the number of experts. MoLE extends the conventional MoE architecture by introducing latent mappings $B_{\text{up}}$, $B_{\text{gate}}$, and $B_{\text{down}}$ that capture shared information across experts. Expert-specific information is encapsulated in the mappings $A_{\text{up}}^{i}$, $A_{\text{down}}^{i}$, and $A_{\text{gate}}^{i}$ for each expert $i$. This latent space decomposition enables MoLE to significantly reduce both memory requirements and computational complexity in FFN layers while maintaining model expressivity.
We introduce Mixture of Latent Experts (MoLE), a novel parameterization paradigm that addresses the core inefficiencies of traditional MoE architectures. Our key insight is that expert modules in standard MoE models contain significant redundancy and operate in unnecessarily high-dimensional spaces. MoLE reformulates each expert operation through a mathematically principled two-phase transformation: (1) a shared projection into a compressed latent space, followed by (2) expert-specific transformations within this lower-dimensional manifold.
Formally, MoLE implements this insight by factorizing each expert’s weight matrix $W^{i}\in\mathbb{R}^{m\times n}$ with $m\ll n$ into the product $A^{i}B$, where $A^{i}\in\mathbb{R}^{m\times m}$ represents expert-specific transformations and $B\in\mathbb{R}^{m\times n}$ represents a shared projection into a latent space of dimension. This factorization yields substantial parameter reduction, particularly in contemporary LLM architectures where the hidden dimension $n$ significantly exceeds the MoE intermediate dimension $m$.
Our contributions are as follows:
  • We propose Mixture of Latent Experts (MoLE), a parameter-efficient architecture that achieves competitive performance with standard MoE models while requiring significantly fewer parameters and reduced computational overhead.
  • We develop a theoretically grounded framework for transforming pre-trained MoE models into the MoLE architecture, including a mathematical characterization of optimal factorization conditions and an efficient two-stage algorithm incorporating low-rank approximation techniques.
  • Through comprehensive empirical evaluation on multiple benchmark datasets, we demonstrate that MoLE preserves or enhances model capabilities across diverse language tasks while substantially improving parameter efficiency, thereby enabling more economical scaling of large language models.
2 Related Works
Finer-Grained Mixture of Experts. Mixture of Experts (MoE), initially introduced by Jacobs et al. (1991) and Jordan and Jacobs (1994), has garnered significant attention in recent years (Aljundi et al., 2017; Collobert et al., 2001; Deisenroth and Ng, 2015; Eigen et al., 2013; Rasmussen and Ghahramani, 2001; Shahbaba and Neal, 2009; Theis and Bethge, 2015). Lepikhin et al. (2020) pioneered the integration of MoE technology into transformer architectures, enabling substantial parameter scaling while maintaining computational efficiency. Subsequently, numerous studies have advanced MoE algorithms, particularly focusing on replacing feed-forward network (FFN) layers with MoE layers in transformer-based Large Language Models (LLMs) (Dai et al., 2024; Du et al., 2022; Fedus et al., 2022; Xue et al., 2024; Zoph et al., 2022).
However, conventional GShard models exhibit limitations in capturing domain-specific expertise due to their relatively small number of experts. To address this constraint and enhance expert specialization, finer-grained MoE architectures were proposed by Dai et al. (2024) and subsequently implemented in several state-of-the-art models (Guo et al., 2025; Liu et al., 2024; Team, 2024). In contrast to traditional GShard MoE designs, finer-grained variants incorporate substantially more experts, each with reduced parameter counts, enabling greater specialization in domain-specific knowledge representation and processing. This approach not only refines the decomposition of knowledge across experts, facilitating more precise learning, but also enhances the flexibility of expert activation combinations, allowing for more specialized and targeted knowledge capture.
Algorithmic Design of MoE. The introduction of expert modules in LLMs introduces several algorithmic challenges that must be addressed during both training and inference phases. A critical aspect of MoE design is the gating function, which orchestrates the engagement of expert computations and the combination of their respective outputs. The gating mechanisms can be broadly categorized into three types: sparse, which activates a subset of experts; dense, which activates all experts; and soft, which encompasses fully-differentiable approaches including input token merging and expert merging (Pan et al., 2024; Zadouri et al., 2023; Puigcerver et al., 2022).
Sparse token-choice gating, where the gating function selects top-k experts for each input token, is the most prevalent approach (Fedus et al., 2022; Lepikhin et al., 2020; Zoph et al., 2022). This method is often augmented with auxiliary loss functions to promote balanced expert utilization (Lepikhin et al., 2020; Fedus et al., 2022; Du et al., 2022). Alternative approaches include expert-choice gating, where each expert selects the top-k tokens they will process (Zhou et al., 2022; 2023), and non-trainable gating mechanisms that use predetermined routing strategies (Roller et al., 2021; Costa et al., 2022; Gururangan et al., 2021).
A promising recent development in MoE research is the integration with parameter-efficient fine-tuning (PEFT) techniques, creating Mixture of Parameter-Efficient Experts (MoPEs) (Zhang et al., 2021; Wu et al., 2022; Ye et al., 2023). These approaches combine the task versatility of MoE with the resource efficiency of PEFT, positioning them as a significant advancement in efficient multi-task learning.
3 MoLE: A Novel Approach to Expert-based Model Architecture
In this paper, we introduce the Mixture of Latent Experts (MoLE), a novel framework that maps experts into latent space to address several limitations of traditional Mixture of Experts (MoE) architectures. For clarity and focus, we exclude shared-experts from our analysis throughout this paper.
3.1 Background: Finer-Grained MoE Architecture
Finer-grained MoE architectures for Feed-Forward Networks (FFNs) employ $N$ distinct experts (Dai et al., 2024). For each expert $E_{i}(x)$ where $i\in\{1,2,\ldots,N\}$, the computation is defined as:
$E_{i}(x):=W^{i}_{\text{down}}\left(W^{i}_{\text{up}}(x)\odot\text{Act}\left(W_{\text{gate}}(x)\right)\right),$ (3.1)
where Act represents the activation function, $W_{\text{up}},W_{\text{gate}}\in\mathbb{R}^{m\times n}$, and $W_{\text{down}}\in\mathbb{R}^{n\times m}$ are linear operators. In this context, $n$ denotes the hidden dimension and $m$ represents the MoE intermediate dimension, where typically $m\leq n$. For input $x$, the FFN layer output is computed as:
$y=x+\sum_{i=1}^{N}g_{i}(x)E_{i}(x),$ (3.2)
where $g_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the router function that determines the contribution of each expert.
While empirical evidence demonstrates that increasing the number of experts leads to superior performance across various applications, this approach introduces significant challenges. The proliferation of parameters results in substantially increased storage requirements and all-to-all network communication overhead, limiting scalability and efficiency.
3.2 Mixture of Latent Experts: Concept and Design
To address these limitations, we propose the Mixture of Latent Experts (MoLE) framework, which fundamentally reimagines the structure of FFN layers in expert-based systems.
Our approach is informed by a careful analysis of the expert computation in Equation equation 3.1, which comprises three distinct operations:
  1. 1.
    Projection in: The input $x$ is mapped from the high-dimensional space $\mathbb{R}^{n}$ to a lower-dimensional space $\mathbb{R}^{m}$ via linear operators $W_{\text{up}}$ and $W_{\text{gate}}$.
  2. 2.
    Non-linear transformation: A one-layer neural network applies a non-linear transformation through the Hadamard product and activation function.
  3. 3.
    Projection out: The intermediate output is mapped back from the low-dimensional space to the original high-dimensional space.
A critical insight from this analysis is that the core functionality of experts primarily stems from the non-linear transformation in the second step. The projection operations in the first and third steps primarily serve to reduce computational complexity, potentially at the cost of limiting the expert’s specialized knowledge capacity.
Drawing inspiration from both Multi-Head Latent Attention (MLA) (Liu et al., 2024), which introduces a “latent space” for KV caches in attention layers, and Grouped-Query Attention (GQA) (Ainslie et al., 2023), which leverages group-based processing, we propose the Mixture of Latent Experts (MoLE). This architecture fundamentally reconsiders how inputs are mapped to a lower-dimensional latent space where expert computation occurs more efficiently.
3.3 Formulation of Mixture of Latent Experts
To formalize the concept of latent experts, we examine the decomposition of expert-specific operators through matrix factorization. Using the “up operator” $W_{\text{up}}^{i}$ of expert $i$ as a representative example, we propose a structured factorization where:
$W_{\text{up}}^{i}=A_{\text{up}}^{i}B_{\text{up}}.$ (3.3)
In this formulation, $B_{\text{up}}\in\mathbb{R}^{m\times n}$ functions as a unified projection operator shared across experts, mapping inputs from the high-dimensional space $\mathbb{R}^{n}$ to a lower-dimensional latent space $\mathbb{R}^{m}$, where typically $m\ll n$. Conversely, $A_{\text{up}}^{i}\in\mathbb{R}^{m\times m}$ represents an expert-specific linear transformation within this latent space, encapsulating the specialized function of expert $i$. Following the terminology established in Multi-Head Latent Attention (MLA), we designate $B_{\text{up}}$ as the latent mapping for the “up operator”.
This factorization principle extends systematically to the “gate operator” as well. On the other side, for the “down operator” $W_{\text{down}}^{i}$, which maps from a lower-dimensional to a higher-dimensional space, the decomposition necessarily assumes a reverse form: $W_{\text{down}}^{i}=B_{\text{down}}A_{\text{down}}^{i}$, where $A_{\text{down}}^{i}\in\mathbb{R}^{m\times m}$ and $B_{\text{down}}\in\mathbb{R}^{n\times m}$.
To optimize the trade-off between model expressivity and parameter efficiency, we introduce a structured grouping mechanism where each subset of $k$ experts shares the same latent mapping matrices $B_{\text{up}}$ and $B_{\text{down}}$. This design establishes a configurable spectrum of architectural possibilities: when $k=1$, each expert maintains its independent latent space, and MoLE becomes functionally equivalent to the standard MoE architecture. Conversely, as $k$ increases, the model achieves progressively higher parameter efficiency at a measured trade-off with expert specialization. This parameterization allows for systematic exploration of the efficiency-performance frontier in mixture-of-experts architectures.
3.4 Formal Definition of MoLE
[…]
$E_{i}(x)=B_{\text{down}}^{\lfloor i/k\rfloor}A_{\text{down}}^{i}\left(A_{\text{up}}^{i}B_{\text{up}}^{\lfloor i/k\rfloor}(x)\odot\text{Act}\left(A_{\text{gate}}^{i}B_{\text{gate}}^{\lfloor i/k\rfloor}(x)\right)\right).$ (3.4)
Consequently, the output of the FFN layer employing the MoLE architecture is computed as:
$y=x+\sum_{i=1}^{N}g_{i}(x)E_{i}(x).$ (3.5)
This formulation effectively disentangles expert-specific computations from the shared dimensionality reduction operations, enabling significant parameter efficiency while preserving model expressivity. The visual comparison between MoE and MoLE architectures is shown in Figure 1.
3.5 Efficiency Benefits of MoLE
The MoLE architecture offers significant efficiency advantages over standard fine-grained MoE models, spanning multiple computational dimensions from memory usage to communication overhead. To quantify these benefits, we provide a comparative analysis between MoE and MoLE for a single FFN layer, assuming identical configurations with hidden dimension $n$, MoE intermediate dimension $m$, and number of experts $N$.
Architecture Parameter Count FLOPs per Forward Pass
Standard MoE $3Nmn$ $(3mn+2m)N$
MoLE (Ours) $3Nm^{2}+3\lfloor\frac{N}{k}\rfloor mn$ $\left(3m^{2}+2m\right)N+3\lfloor\frac{N}{k}\rfloor mn$
Table 1: Efficiency comparison between standard MoE and our proposed MoLE architectures for a single FFN layer. Parameter count represents the total number of trainable parameters in expert modules, while FLOPs measures the computational cost during a forward pass with top-1 routing. Here, $n$ represents the hidden dimension, $m$ the MoE intermediate dimension, $N$ the number of experts, and $k$ the number of latent experts per group.
Parameter Efficiency As shown in Table 1, MoLE substantially reduces the parameter count compared to standard MoE, particularly when $m\ll n$, which is the typical case in modern LLMs. For instance, in DeepSeek-V3, $n=7168$ while $m=2048$. The parameter reduction stems from our latent parameterization, where expert-specific transformations operate in the lower-dimensional latent space ($m\times m$) rather than directly on the high-dimensional hidden space ($m\times n$).
Computational Efficiency Beyond parameter savings, MoLE also reduces the computational cost measured in FLOPs. The efficiency gain becomes particularly pronounced when the number of experts $N$ is large and $k$ is small (meaning fewer latent projection matrices are used). This computational advantage translates to faster inference and training times, especially on hardware where memory bandwidth is a bottleneck.
Communication Overhead Reduction A critical but often overlooked benefit of MoLE is the reduction in all-to-all communication costs during distributed training and inference. In standard MoE models, the full expert parameters ($3Nmn$ in total) must be synchronized across devices. In contrast, MoLE requires synchronization of significantly fewer parameters, reducing network bandwidth requirements and improving scalability for distributed deployments.
Memory Access Patterns MoLE also offers improved cache efficiency during computation. The smaller matrices used in latent transformations ($A^{i}\in\mathbb{R}^{m\times m}$) exhibit better locality of reference compared to the larger matrices in standard MoE ($W^{i}\in\mathbb{R}^{m\times n}$), potentially leading to higher utilization of fast cache memory and reduced main memory bandwidth demands.
4 Theoretical Analysis of MoE to MoLE Transformation
In this section, we establish the theoretical foundation for transforming a standard Mixture of Experts (MoE) model into its corresponding Mixture of Latent Experts (MoLE) counterpart. We focus on the “up operator” as a representative example. For analytical clarity, we make two simplifying assumptions: (1) we omit the subscript of $W_{\text{up}}^{i}$ for ease of notation during this analysis, and (2) we consider the case where $k=N$, implying a single shared latent space operator for all experts.
Keep the “down operator”. Note that for the “down operator” $W_{\text{down}}$, it maps the intermediate term from a low dimension into a high dimension, which captures the intrinsic properties of the given expert. Moreover, empirical experiments in Appendix C show that maintaining the structure of the “down operator” performs much better than not. Hence, we maintain the “down operator” in the transformed MoLE architecture and only change the “up operator” and “gate operator” from the MoE structure into the MoLE structure.
4.1 Transformation via Matrix Factorization
For a given weight matrix $W^{i}$ associated with expert $i$, we aim to find corresponding matrices $A^{i}$ and $B$ such that $W^{i}x\approx A^{i}Bx$ for any input $x$. This transformation represents the post-training perspective where we seek to convert pre-trained MoE parameters into the MoLE architecture.
One direct approach is to determine matrices $A^{i}$ and $B$ such that $A^{i}B\approx W^{i}$ for all $i\in\{1,2,\ldots,N\}$. This leads naturally to the following optimization problem:
$\min_{A^{i},B}\quad F(A^{1},\cdots,A^{N},B):=\frac{1}{2}\sum_{i=1}^{N}\|W^{i}-A^{i}B\|_{F}^{2},$ (4.1)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. Problem equation 4.1 admits the theoretical optimal solutions by using the SVD decomposition technique.
Closed-Form Solution via Singular Value Decomposition. Obviously, problem equation 4.1 has infinitely many solutions since one can change one optimal solution $(A^{*},B^{*})$ into another by selecting a constant $\lambda$ and obtain another optimal solution $(\lambda A^{*},\frac{1}{\lambda}B^{*})$. Hence, we only provide “one” optimal solution here.
To derive a closed-form solution to problem equation 4.1, we consolidate the matrices into the following block structures:
$W=\begin{pmatrix}W^{1}\\ W^{2}\\ \vdots\\ W^{N}\end{pmatrix},\quad A=\begin{pmatrix}A^{1}\\ A^{2}\\ \vdots\\ A^{N}\end{pmatrix}.$ (4.2)
With this notation, problem equation 4.1 can be reformulated as:
$\min_{A,B}\quad\frac{1}{2}\|W-AB\|_{F}^{2},$ (4.3)
where $A\in\mathbb{R}^{mN\times m}$ and $B\in\mathbb{R}^{m\times n}$.
While problem equation 4.3 has infinitely many solutions due to its underdetermined nature, the Eckart-Young-Mirsky theorem (Schmidt, 1989) provides an optimal solution with respect to the Frobenius norm. Specifically, let $W=U\Sigma V^{\top}$ be the singular value decomposition (SVD) of $W$, where:
  • $U\in\mathbb{R}^{mN\times mN}$ is an orthogonal matrix whose columns are the left singular vectors of $W$
  • $\Sigma\in\mathbb{R}^{mN\times n}$ is a rectangular diagonal matrix with singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{\min(n,mN)}\geq 0$ on its diagonal
  • $V\in\mathbb{R}^{n\times n}$ is an orthogonal matrix whose columns are the right singular vectors of $W$
Let $\Sigma_{m}$ be the truncated version of $\Sigma$ that retains only the $m$ largest singular values: $\Sigma_{m}=\text{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{m},0,\ldots,0)\in\mathbb{R}^{mN\times n}$. According to the Eckart-Young-Mirsky theorem, an optimal solution to problem equation 4.3 is given by:
$A^{*}=U_{m}\Sigma_{m}^{1/2},\quad B^{*}=\Sigma_{m}^{1/2}V_{m}^{\top},$ (4.4)
where $U_{m}\in\mathbb{R}^{mN\times m}$ consists of the first $m$ columns of $U$, $\Sigma_{m}^{1/2}\in\mathbb{R}^{m\times m}$ is a diagonal matrix containing the square roots of the $m$ largest singular values, and $V_{m}\in\mathbb{R}^{n\times m}$ consists of the first $m$ columns of $V$. This factorization yields the minimum Frobenius norm error among all rank-$m$ approximations of $W$, with the approximation error given by:
$\|W-A^{*}B^{*}\|_{F}^{2}=\sum_{i=m+1}^{\min(n,mN)}\sigma_{i}^{2}.$ (4.5)
4.2 Minimizing Factorization Residuals
In the previous subsection, we established methods for transforming MoE models into their MoLE counterparts from a theoretical perspective. Here, we address a critical aspect of this transformation: minimizing the residual error that inevitably arises when factorizing expert weights. We begin by establishing the precise conditions under which exact factorization is possible.
Theorem 1.
Given matrices $W^{i}\in\mathbb{R}^{m\times n}$ with $m\leq n$, there exist matrices $A^{i}\in\mathbb{R}^{m\times m}$ and a common matrix $B\in\mathbb{R}^{m\times n}$ such that $A^{i}B=W^{i}$ for all $i\in\{1,2,\ldots,N\}$, if and only if there exists an $(n-m)$-dimensional subspace $K\subseteq\mathbb{R}^{n}$ satisfying:
$K\subseteq\bigcap_{i=1}^{N}\ker(W^{i}).$ (4.6)
Proof.
Necessity: Suppose there exist $B\in\mathbb{R}^{m\times n}$ and $A^{i}\in\mathbb{R}^{m\times m}$ such that $W^{i}=A^{i}B$ for all $i$. Since $m\leq n$ and we require exact factorization, $B$ must be row-full-rank (i.e., $\text{rank}(B)=m$). Consequently, its right nullspace $\ker(B)$ has dimension $n-m$. For any vector $x\in\ker(B)$, we have:
$W^{i}x=A^{i}Bx=A^{i}\cdot 0=0,$ (4.7)
which implies $\ker(B)\subseteq\ker(W^{i})$ for all $i$. Setting $K=\ker(B)$, we obtain an $(n-m)$-dimensional subspace contained in the intersection of all $\ker(W^{i})$.
Sufficiency: Suppose there exists an $(n-m)$-dimensional subspace $K\subseteq\mathbb{R}^{n}$ such that $K\subseteq\ker(W^{i})$ for all $i$. We can construct $B\in\mathbb{R}^{m\times n}$ such that $\ker(B)=K$. Since $\dim(K)=n-m$, the matrix $B$ has rank $m$ by the rank-nullity theorem. For each $W^{i}$, the inclusion $K\subseteq\ker(W^{i})$ implies that any vector mapped to zero by $B$ is also mapped to zero by $W^{i}$. By the fundamental theorem of linear algebra, this means $\text{Row}(W^{i})\subseteq\text{Row}(B)$, where $\text{Row}(\cdot)$ denotes the row space. Therefore, there exists $A^{i}\in\mathbb{R}^{m\times m}$ such that $W^{i}=A^{i}B$ for each $i$. ∎
Theorem 1 provides a geometric interpretation of the factorization problem: exact factorization is possible only when the nullspaces of all expert matrices share a sufficiently large common subspace. In practical LLM implementations, however, this condition is rarely satisfied for FFN layers in MoE models, as our empirical analysis confirms.
Given that exact factorization is generally unattainable, we now consider how to minimize the approximation error through strategic rank reduction. The rank-nullity theorem (Lang, 1987) states that for any linear mapping $W:X\rightarrow Y$:
$\text{rank}(W)+\dim(\ker(W))=\dim(X).$ (4.8)
In our context, $X$ represents the hidden space ($\mathbb{R}^{n}$) and $Y$ the MoE intermediate space ($\mathbb{R}^{m}$). Therefore, $\text{rank}(W^{i})+\dim(\ker(W^{i}))=n$ for all $i\in\{1,2,\ldots,N\}$.
This relationship suggests a strategic approach: by reducing the rank of each $W^{i}$, we can increase the dimension of its nullspace. Specifically, if we constrain each $W^{i}$ to have rank $r<m$, then $\dim(\ker(W^{i}))=n-r>n-m$. This increases the probability of finding a substantial common subspace within the intersection $\bigcap_{i=1}^{N}\ker(W^{i})$, thereby improving the quality of our factorization.
We implement this approach by computing low-rank approximations of each $W^{i}$ before attempting factorization. Importantly, our empirical experiments in Appendix A demonstrate that this rank reduction has minimal impact on model performance, suggesting that these FFN operators in MoE models inherently possess low-rank structure that can be exploited for more efficient parameterization.
4.3 Transfer MoE to MoLE: A Unified Framework
Based on our theoretical analyses in Sections 4.1 and 4.2, we now propose a unified, systematic framework for transforming Mixture of Experts (MoE) models into their more parameter-efficient Mixture of Latent Experts (MoLE) counterparts. Our framework consists of two principal steps, carefully designed to preserve model capabilities while enabling the latent parameterization.
Step 1: Rank Reduction. For each expert operator $W^{i}$, we compute a low-rank approximation $\tilde{W}^{i}$ that maintains the essential functionality of the original operator while increasing the dimension of its nullspace. This step is motivated by our theoretical analysis showing that larger nullspace intersections facilitate better factorization. We determine the optimal rank based on empirical validation to ensure minimal performance degradation.
Step 2: Matrix Factorization. Using the rank-reduced operators $\{\tilde{W}^{i}\}_{i=1}^{N}$, we apply matrix factorization techniques to identify the shared projection matrix $B$ and the expert-specific latent transformations $\{A^{i}\}_{i=1}^{N}$. For this step, we employ the SVD-based approach detailed in Section 4.1, which provides the optimal factorization with respect to the Frobenius norm.
We formalize our approach in Algorithm 1, which provides a complete computational procedure for transforming MoE parameters into the MoLE architecture.
Algorithm 1 Transformation of MoE to MoLE
1:Expert weight matrices $\{W^{i}\}_{i=1}^{N}$, target rank $r$, latent dimension $m$
2:Latent expert matrices $\{A^{i}\}_{i=1}^{N}$, shared projection matrix $B$
3:// Step 1: Rank Reduction
4:for $i=1$ to $N$ do
5:     Compute SVD: $W^{i}=U^{i}\Sigma^{i}(V^{i})^{\top}$
6:     Truncate to rank $r$: $\tilde{W}^{i}=U^{i}[:,:r]\cdot\Sigma^{i}[:r,:r]\cdot(V^{i}[:,:r])^{\top}$
7:end for
8:// Step 2: Matrix Factorization
9:Construct concatenated matrix: $\tilde{W}=[\tilde{W}^{1};\tilde{W}^{2};\ldots;\tilde{W}^{N}]$
10:Compute SVD: $\tilde{W}=U\Sigma V^{\top}$
11:Extract first $m$ singular values and vectors
$A=U[:,:m]\cdot\Sigma[:m,:m]^{1/2}$
[…]
14:Partition $A$ into $N$ blocks to obtain $\{A^{i}\}_{i=1}^{N}$
15:return $\{A^{i}\}_{i=1}^{N},B$
This algorithm provides a computationally efficient procedure for transforming standard MoE layers into MoLE architecture. The rank reduction parameter $r$ and the latent dimension $m$ serve as hyperparameters that can be tuned to balance performance preservation against parameter efficiency. Typically, we set $r\leq m<\min(n,\sum_{i=1}^{N}m)$, where $n$ is the input dimension.
Our framework is applicable to all linear operators within an MoE layer, including the up, down, and gate operators. By applying this transformation to each operator independently, we convert the entire MoE model to MoLE while maintaining model quality, as demonstrated in our experimental evaluations.
5 Experiments
This section presents a comprehensive comparative analysis between Mixture of Latent Experts (MoLE) and Mixture of Experts (MoE) architectures in the context of pretraining GPT-2 (Radford et al., 2019) models and evaluating their subsequent performance on downstream tasks.
5.1 Experimental Setup
The MoE and MoLE models were configured with identical architectural parameters except for their Feed-Forward Network (FFN) layer structures, as detailed in Table 6. Both architectures implement $N=32$ experts in their respective FFN layers. For the MoLE model, we establish $k=8$, indicating that each group of $8$ experts shares a common latent representation space. All models were trained on the Wikipedia English (Foundation, ) with sequences truncated to a maximum length of $512$ tokens. Training was conducted using the AdamW optimizer with consistent hyperparameters across all experimental conditions. The comprehensive specifications of model architectures and training configurations are presented in Table 6.
5.2 Parameter Efficiency
As quantified in Table 6, the introduction of shared latent spaces in MoLE architecture yields a substantial reduction in model parameter count compared to the standard MoE architecture. Specifically, MoLE achieves a 40% reduction in non-embedding parameters while maintaining comparable representational capacity. This parameter efficiency represents a significant advancement in model scalability without sacrificing performance.
5.3 Training Dynamics
Figure 2 illustrates the pretraining convergence characteristics of both MoE and MoLE models. The training loss trajectories reveal that MoLE maintains competitive optimization dynamics despite its significantly reduced parameter count. Although the MoE model converges to marginally lower loss values, this difference is negligible when considering the substantial parameter efficiency gained with MoLE. These results suggest that the shared latent representation in MoLE effectively preserves the essential modeling capacity while eliminating redundant parameterization inherent in traditional MoE architectures.
Figure 2: Comparison of training loss curves between MoE and MoLE models on the English Wikipedia dataset. The convergence patterns demonstrate comparable performance, with MoLE exhibiting slightly higher loss values while utilizing substantially fewer parameters than the standard MoE architecture.
6 Conclusion
In this paper, we presented Mixture of Latent Experts (MoLE), a novel parameter-efficient architecture that addresses the fundamental limitations of traditional MoE models. By factorizing expert weight matrices into shared projections and expert-specific transformations within a lower-dimensional latent space, MoLE significantly reduces parameter count and computational overhead without sacrificing model performance. Our comprehensive empirical evaluation demonstrates that MoLE not only preserves the capabilities of standard MoE models across diverse language tasks but also enables more economical scaling of large language models.
The theoretical framework we developed for transforming pre-trained MoE models into the MoLE architecture provides a principled approach to expert parameterization, offering insights into the nature of redundancy in neural networks. As models continue to grow in scale and complexity, architectures like MoLE that prioritize parameter efficiency while maintaining performance will become increasingly valuable for practical deployment scenarios. Future work could explore extending these factorization techniques to other components of the transformer architecture and investigating dynamic adaptation of the latent space dimensionality based on input complexity.
Appendix A Parameter Redundancy in Neural Networks
This section provides empirical evidence for significant parameter redundancy within Feed-Forward Network (FFN) layers, thereby substantiating the theoretical framework presented in Section 4. We conduct our analysis using Qwen1.5-MoE-A2.7B [Team, 2024], a state-of-the-art Mixture-of-Experts architecture comprising $14.3$ billion total parameters while activating only $2.7$ billion parameters during inference. Each FFN layer in this architecture is characterized by an intermediate dimension $m=1408$, hidden dimension $n=2048$, and $N=60$ distinct expert modules.
Figure 3: Singular value distributions of expert modules in the Qwen1.5-MoE-A2.7B model. The relatively uniform distributions indicate that weight matrices exhibit high-rank characteristics, with no significant concentration of information in a small subset of singular values.
As illustrated in Figure 3, the singular value distributions of the expert modules demonstrate an absence of dominant singular values—a finding that contradicts conventional assumptions about the inherent low-rank structure of neural network weight matrices. This observation suggests that naive dimensional reduction techniques applying uniform low-rank approximations would potentially eliminate performance-critical parameters.
Nevertheless, we demonstrate that despite the apparent full-rank nature of the weight matrices $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$, this characteristic is not a prerequisite for optimal model performance. Through systematic experimentation, we establish that these operators can be effectively replaced with carefully calibrated low-rank approximations while maintaining comparable performance metrics.
For our analysis, we define a ratio-$r$ low-rank approximation $\tilde{W}$ of any matrix $W$ as a matrix whose rank equals $r\times\text{rank}(W)$, where $0<r\leq 1$. In accordance with the Eckart-Young-Mirsky theorem [Schmidt, 1989], these approximations are computed via Singular Value Decomposition (SVD), retaining only the largest $r\times\text{rank}(W)$ singular values and their corresponding singular vectors.
To rigorously assess model capabilities under low-rank constraints, we evaluate performance on three benchmark tasks: MMLU [Hendrycks et al., 2021], GSM8K [Cobbe et al., 2021], and Wikitext-2 [Merity et al., 2016]—the same evaluation suite utilized in the original Qwen report. All experiments were conducted using the lm-evaluation-harness evaluation framework [Gao et al., 2024]. Table 2 presents the comparative results.
Low-rank ratio $r$ GSM8K (%) $\uparrow$ MMLU (%) $\uparrow$ Wikitext PPL $\downarrow$
$1.0$ (Original) 60.1 61.0 9.49
$0.8$ 61.2 60.3 9.65
$0.6$ 60.1 59.4 9.85
Table 2: Performance comparison of QwenMoE models with varying low-rank approximation ratios across multiple benchmarks. For the low-rank ratio $r=0.8$ case, the low-rank estimation model performs competitive with the original model on various tasks.
Table 2 illustrates the relationship between rank reduction and model performance. The baseline case ($r=1.0$) represents the original, unmodified model with full-rank weight matrices. Notably, when reducing the rank of FFN operators by 20% ($r=0.8$), we observe no significant performance degradation. In fact, the reduced-rank model demonstrates superior performance on the GSM8K benchmark, achieving a 1.1 percentage point improvement over the original model, while maintaining comparable performance on MMLU accuracy and Wikitext-2 PPL.
These empirical findings provide compelling evidence that, despite the mathematical full-rank property of FFN weight matrices, a substantial proportion of parameters contain redundant information that can be effectively approximated through lower-dimensional representations. This parameter redundancy phenomenon forms the empirical foundation for our theoretical analysis in Section 4.
Appendix B Transformation from MoE to MoLE
This section presents our empirical analysis of transforming standard MoE architectures into their corresponding MoLE counterparts. We specifically investigate the Qwen1.5-MoE-A2.7B model and systematically examine the impact of varying the latent space configuration parameter $k$ on model performance.
As established in Section 3, the parameter $k$ in MoLE architecture represents a critical hyperparameter that controls the trade-off between parameter efficiency and model expressivity. We evaluate three distinct configurations of $k\in\{1,10,60\}$, where:
  • $k=1$ corresponds to the original MoE architecture (baseline)
  • $k=10$ represents a balanced MoLE configuration with multiple latent spaces
  • $k=60$ represents the extreme case with a single shared latent space
For consistent evaluation with Section A, we assess model performance on three established benchmarks: MMLU (reasoning), GSM8K (mathematical reasoning), and Wikitext-2 (language modeling perplexity).
Table 3 presents the comparative performance across different MoLE configurations.
Number of experts per latent ($k$) GSM8K (%) $\uparrow$ MMLU (%) $\uparrow$ Wikitext-2 PPL $\downarrow$
$1$ (Original MoE) 60.1 61.0 9.49
$10$ 59.4 60.0 9.74
$60$ 14.3 40.1 25.97
Table 3: Performance comparison of Qwen1.5-MoE-A2.7B with varying MoLE configurations. The balanced configuration ($k=10$) maintains near-parity with the original model while achieving significant parameter reduction. The single latent space configuration ($k=60$) demonstrates substantial performance degradation.
The results demonstrate a clear relationship between latent space configuration and model performance. The balanced MoLE configuration ($k=10$) achieves comparable performance to the original MoE architecture, with only marginal degradation across all benchmarks: a 0.7 percentage point decrease on GSM8K, 1.0 percentage point decrease on MMLU, and 2.6% increase in perplexity on Wikitext-2. This configuration successfully reduces parameter count while preserving most of the model’s capabilities.
In contrast, the extreme configuration ($k=60$) with a single shared latent space per FFN layer exhibits catastrophic performance degradation: a 45.8 percentage point decrease on GSM8K, 20.9 percentage point decrease on MMLU, and 173.7% increase in perplexity. This dramatic performance collapse empirically validates our theoretical analysis that multiple latent spaces are necessary to preserve the expressive capacity of the original MoE architecture.
These findings provide compelling evidence that the MoLE architecture, when properly configured with an appropriate number of latent spaces, can achieve substantial parameter efficiency while maintaining competitive performance. Furthermore, they establish clear boundaries for the practical application of latent space factorization in expert-based neural architectures.
Appendix C Critical Role of the “Down Operator”
This section presents empirical evidence demonstrating that the “down operator” encapsulates more essential information than other components in Mixture of Experts (MoE) Large Language Models (LLMs). Through systematic experimentation, we establish the importance of preserving this operator’s structure for maintaining model performance.
Consistent with our methodology in Appendix B, we utilize the Qwen1.5-MoE-A2.7B model with a fixed latent parameter $k=10$. To isolate the significance of the “down operator,” we implement two distinct transformation approaches:
  1. 1.
    Partial transformation (“up+gate”) - Converts only the “up operator” and “gate operator” to their Mixture of Latent Experts (MoLE) equivalents while preserving the original “down operator.”
  2. 2.
    Complete transformation (“all”) - Transforms all three components (“up,” “gate,” and “down” operators) into their corresponding latent space representations.
We evaluate these transformations across multiple downstream tasks, with results summarized in Table 4.
Transformation Type ($k=10$) GSM8K (%) $\uparrow$ MMLU (%) $\uparrow$ Wikitext-2 PPL $\downarrow$
Original MoE 60.1 61.0 9.49
Partial (“up+gate”) 59.4 60.0 9.74
Complete (“all”) 43.2 57.0 10.63
Table 4: Performance comparison of Qwen1.5-MoE-A2.7B under different transformation configurations. Results demonstrate the critical importance of preserving the “down operator” structure for maintaining model performance.
The results in Table 4 reveal a clear performance hierarchy. The original MoE model achieves optimal performance across all metrics. The partial transformation, which preserves the “down operator,” exhibits only marginal performance degradation (0.7% decrease on GSM8K, 1.0% on MMLU, and 0.25 increase in perplexity). In stark contrast, the complete transformation that modifies all operators, including the “down operator,” suffers substantial performance deterioration (16.9% decrease on GSM8K, 4.0% on MMLU, and 1.14 increase in perplexity).
These findings provide compelling evidence that the “down operator” contains critical information that significantly impacts model performance. When this operator is transformed into the latent space, substantial information loss occurs, resulting in markedly diminished capabilities across reasoning and knowledge-intensive tasks. This asymmetric importance among the operators suggests that architectural modifications to MoE models should prioritize preserving the structure of the “down operator” to maintain performance integrity.
Appendix D Downstream Task Evaluation
To rigorously assess the generalization capabilities of the proposed MoLE architecture, we conducted comprehensive evaluations across a diverse set of downstream tasks. The comparative performance metrics are presented in Table 5.
Model Wikitext-2
(PPL $\downarrow$)
Standard MoE 75.86
MoLE 81.57
Table 5: Performance comparison between standard MoE and MoLE models across language modeling benchmark tasks. Bold values indicate superior performance.
The empirical results in Table 5 demonstrate that despite utilizing approximately 40% fewer parameters than the standard MoE architecture, the proposed MoLE model achieves comparable performance across the PPL tasks. This parameter efficiency without substantial performance degradation highlights the effectiveness of our proposed architectural modification.
Appendix E Training Arguments
Hyperparameters MoE MoLE
FFN layers size $151$M $94$M
Vocabulary size $50257$ $50257$
Number of layers $12$ $12$
Number of attention heads $8$ $8$
Hidden dimension $n$ $512$ $512$
Intermediate dimension $1024$ $1024$
MoE intermediate dimension $m$ $256$ $256$
Number of experts $N$ $32$ $32$
Experts per latent space $k$ $1$ $8$
Load balancing mechanism Auxiliary loss Auxiliary loss
Optimizer AdamW AdamW
Learning rate $3\times 10^{-4}$ $3\times 10^{-4}$
Learning rate schedule Cosine decay Cosine decay
Table 6: Model architecture and training hyperparameter configurations for MoE and MoLE models.

### v2

@@
-Beyond Standard MoE: Mixture of Latent Experts for Resource-Efficient Language Models
+MoLAE: Mixture of Latent Experts for Parameter-Efficient Language Models
 Abstract
-Mixture of Experts (MoE) has emerged as a pivotal architectural paradigm for efficient scaling of Large Language Models (LLMs), operating through selective activation of parameter subsets for each input token.
-Nevertheless, conventional MoE architectures encounter substantial challenges, including excessive memory utilization and communication overhead during training and inference, primarily attributable to the proliferation of expert modules.
-In this paper, we introduce Mixture of Latent Experts (MoLE), a novel parameterization methodology that facilitates the mapping of specific experts into a shared latent space.
-Specifically, all expert operations are systematically decomposed into two principal components: a shared projection into a lower-dimensional latent space, followed by expert-specific transformations with significantly reduced parametric complexity.
-This factorized approach substantially diminishes parameter count and computational requirements.
-Beyond the pretraining implementation of the MoLE architecture, we also establish a rigorous mathematical framework for transforming pre-trained MoE models into the MoLE architecture, characterizing the sufficient conditions for optimal factorization and developing a systematic two-phase algorithm for this conversion process.
-Our comprehensive theoretical analysis demonstrates that MoLE significantly enhances computational efficiency across multiple dimensions while preserving model representational capacity.
-Empirical evaluations corroborate our theoretical findings, confirming that MoLE achieves performance comparable to standard MoE implementations while substantially reducing resource requirements.
+Mixture of Experts (MoE) has become a key architectural paradigm for efficiently scaling Large Language Models (LLMs) by selectively activating a subset of parameters for each input token.
+However, standard MoE architectures face significant challenges, including high memory consumption and communication overhead during distributed training.
+In this paper, we introduce Mixture of Latent Experts (MoLAE), a novel parameterization that addresses these limitations by reformulating expert operations through a shared projection into a lower-dimensional latent space, followed by expert-specific transformations.
+This factorized approach substantially reduces parameter count and computational requirements, particularly in existing LLMs where hidden dimensions significantly exceed MoE intermediate dimensions.
+We provide a rigorous mathematical framework for transforming pre-trained MoE models into MoLAE architecture, characterizing conditions for optimal factorization, and developing a systematic two-step algorithm for this conversion.
+Our comprehensive theoretical analysis demonstrates that MoLAE significantly improves efficiency across multiple dimensions while preserving model capabilities.
+Experimental results confirm that MoLAE achieves comparable performance to standard MoE with substantially reduced resource requirements.
 1 Introduction
@@
 This empirical observation motivates a fundamental rethinking of expert parameterization in neural architectures.
-Figure 1: Architectural comparison between MoE and MoLE in the feed-forward network (FFN) layer.
-In both diagrams, $N$ denotes the number of experts.
-MoLE extends the conventional MoE architecture by introducing latent mappings $B_{\text{up}}$, $B_{\text{gate}}$, and $B_{\text{down}}$ that capture shared information across experts.
+Figure 1: Architectural comparison between MoE and MoLAE in the FFN layer.
+In both diagrams, $N_{r}$ denotes the number of routed experts.
+MoLAE extends the conventional MoE architecture by introducing latent mappings $B_{\text{up}}$, $B_{\text{gate}}$, and $B_{\text{down}}$ that capture shared information across experts.
 Expert-specific information is encapsulated in the mappings $A_{\text{up}}^{i}$, $A_{\text{down}}^{i}$, and $A_{\text{gate}}^{i}$ for each expert $i$.
-This latent space decomposition enables MoLE to significantly reduce both memory requirements and computational complexity in FFN layers while maintaining model expressivity.
-We introduce Mixture of Latent Experts (MoLE), a novel parameterization paradigm that addresses the core inefficiencies of traditional MoE architectures.
+In this work, we introduce Mixture of LAtent Experts (MoLAE), a novel parameterization paradigm that addresses the core inefficiencies of traditional MoE architectures.
 Our key insight is that expert modules in standard MoE models contain significant redundancy and operate in unnecessarily high-dimensional spaces.
-MoLE reformulates each expert operation through a mathematically principled two-phase transformation: (1) a shared projection into a compressed latent space, followed by (2) expert-specific transformations within this lower-dimensional manifold.
-Formally, MoLE implements this insight by factorizing each expert’s weight matrix $W^{i}\in\mathbb{R}^{m\times n}$ with $m\ll n$ into the product $A^{i}B$, where $A^{i}\in\mathbb{R}^{m\times m}$ represents expert-specific transformations and $B\in\mathbb{R}^{m\times n}$ represents a shared projection into a latent space of dimension.
+MoLAE reformulates each expert operation through a mathematically principled two-phase transformation: (1) a shared projection into a compressed latent space, followed by (2) expert-specific transformations within this lower-dimensional manifold.
+Formally, MoLAE implements this insight by factorizing each expert’s weight matrix $W^{i}\in\mathbb{R}^{m\times n}$ into the product $A^{i}B$, where $A^{i}\in\mathbb{R}^{m\times m}$ represents expert-specific transformations and $B\in\mathbb{R}^{m\times n}$ represents a shared projection into a latent space of dimension.
 This factorization yields substantial parameter reduction, particularly in contemporary LLM architectures where the hidden dimension $n$ significantly exceeds the MoE intermediate dimension $m$.
-Our contributions are as follows:
-  • We propose Mixture of Latent Experts (MoLE), a parameter-efficient architecture that achieves competitive performance with standard MoE models while requiring significantly fewer parameters and reduced computational overhead.
-  • We develop a theoretically grounded framework for transforming pre-trained MoE models into the MoLE architecture, including a mathematical characterization of optimal factorization conditions and an efficient two-stage algorithm incorporating low-rank approximation techniques.
-  • Through comprehensive empirical evaluation on multiple benchmark datasets, we demonstrate that MoLE preserves or enhances model capabilities across diverse language tasks while substantially improving parameter efficiency, thereby enabling more economical scaling of large language models.
+Our contributions are as follows: 1) We propose MoLAE, a parameter-efficient architecture that achieves competitive performance with standard MoE models while requiring significantly fewer parameters and reduced computational overhead.
+2) We develop a theoretically grounded framework for transforming pre-trained MoE models into the MoLAE architecture, including a mathematical characterization of optimal factorization conditions and an efficient two-stage algorithm incorporating low-rank approximation techniques.
+3) Through comprehensive empirical evaluation on multiple benchmark datasets, we demonstrate that MoLAE preserves or enhances model capabilities across diverse language tasks while substantially improving parameter efficiency, thereby enabling more economical scaling of large language models.
 2 Related Works
@@
 This method is often augmented with auxiliary loss functions to promote balanced expert utilization (Lepikhin et al., 2020; Fedus et al., 2022; Du et al., 2022).
-Alternative approaches include expert-choice gating, where each expert selects the top-k tokens they will process (Zhou et al., 2022; 2023), and non-trainable gating mechanisms that use predetermined routing strategies (Roller et al., 2021; Costa et al., 2022; Gururangan et al., 2021).
-A promising recent development in MoE research is the integration with parameter-efficient fine-tuning (PEFT) techniques, creating Mixture of Parameter-Efficient Experts (MoPEs) (Zhang et al., 2021; Wu et al., 2022; Ye et al., 2023).
+Alternative approaches include expert-choice gating, where each expert selects the top-k tokens they will process (Zhou et al., 2022, 2023), and non-trainable gating mechanisms that use predetermined routing strategies (Roller et al., 2021; Costa et al., 2022; Gururangan et al., 2021).
+A promising recent development in MoE is the integration with parameter-efficient fine-tuning (PEFT) techniques, creating Mixture of Parameter-Efficient Experts (MoPEs) (Zhang et al., 2021; Wu et al., 2022; Ye et al., 2023).
 These approaches combine the task versatility of MoE with the resource efficiency of PEFT, positioning them as a significant advancement in efficient multi-task learning.
-3 MoLE: A Novel Approach to Expert-based Model Architecture
-In this paper, we introduce the Mixture of Latent Experts (MoLE), a novel framework that maps experts into latent space to address several limitations of traditional Mixture of Experts (MoE) architectures.
-For clarity and focus, we exclude shared-experts from our analysis throughout this paper.
-3.1 Background: Finer-Grained MoE Architecture
-Finer-grained MoE architectures for Feed-Forward Networks (FFNs) employ $N$ distinct experts (Dai et al., 2024). For each expert $E_{i}(x)$ where $i\in\{1,2,\ldots,N\}$, the computation is defined as:
-$E_{i}(x):=W^{i}_{\text{down}}\left(W^{i}_{\text{up}}(x)\odot\text{Act}\left(W_{\text{gate}}(x)\right)\right),$ (3.1)
+3 Redundancy in Standard MoE models
+3.1 Background: Standard MoE Architecture
+Finer-grained MoE architectures for FFNs employ $N$ distinct experts (Dai et al., 2024). For each expert $E_{i}(x)$ where $i\in\{1,2,\ldots,N\}$, the computation is defined as:
+$E_{i}(x):=W^{i}_{\text{down}}\left(W^{i}_{\text{up}}(x)\odot\textsc{Act}\left(W_{\text{gate}}(x)\right)\right),$ (1)
 where Act represents the activation function, $W_{\text{up}},W_{\text{gate}}\in\mathbb{R}^{m\times n}$, and $W_{\text{down}}\in\mathbb{R}^{n\times m}$ are linear operators.
@@
 For input $x$, the FFN layer output is computed as:
-$y=x+\sum_{i=1}^{N}g_{i}(x)E_{i}(x),$ (3.2)
+$y=x+\sum_{i=1}^{N}g_{i}(x)E_{i}(x),$ (2)
 where $g_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the router function that determines the contribution of each expert.
@@
 The proliferation of parameters results in substantially increased storage requirements and all-to-all network communication overhead, limiting scalability and efficiency.
+3.2 Parameter Redundancy in MoE Models
+In this section, we provide empirical evidence for significant parameter redundancy within FFN layers, substantiating the theoretical framework presented in Section 5.
+We conduct analysis using the Qwen1.5-MoE-A2.7B (Team, 2024) model, a popular MoE model comprising 14.3B parameters while activating only 2.7B parameters during inference, with 60 distinct experts.
+For our analysis, we define a ratio-$r$ low-rank approximation $\tilde{W}$ of any matrix $W$ as a matrix whose rank equals $r\times\text{rank}(W)$, where $0<r\leq 1$.
+In accordance with the Eckart-Young-Mirsky theorem (Schmidt, 1989), these approximations are computed via Singular Value Decomposition (SVD), retaining only the largest $r\times\text{rank}(W)$ singular values and their corresponding singular vectors.
+To rigorously assess model capabilities under low-rank constraints, we evaluate performance on three benchmark tasks: MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), and Wikitext-2 (Merity et al., 2016).
+All experiments were conducted using the lm-evaluation-harness evaluation framework (Gao et al., 2024).
+Table 1 presents the comparative results.
+Table 1: Performance comparison of MoE models with varying low-rank approximation ratios across multiple benchmarks.
+Low-rank ratio $r$ GSM8K (%) $\uparrow$ MMLU (%) $\uparrow$ Wikitext PPL $\downarrow$
+$1.0$ (Original) 60.1 61.0 9.49
+$0.8$ 61.2 60.3 9.65
+$0.6$ 60.1 59.4 9.85
+Table 1 illustrates the relationship between rank reduction and model performance.
+The baseline case ($r=1.0$) represents the original, unmodified model with full-rank weight matrices.
+Notably, when reducing the rank of FFN operators by 20% ($r=0.8$), we observe no significant performance degradation.
+In fact, the reduced-rank model demonstrates superior performance on the GSM8K benchmark, achieving a 1.1 percentage point improvement over the original model, while maintaining comparable performance on MMLU accuracy and Wikitext-2 PPL.
+These empirical findings provide compelling evidence that, despite the mathematical full-rank property of FFN weight matrices, a substantial proportion of parameters contain redundant information that can be effectively approximated through lower-dimensional representations.
+This parameter redundancy phenomenon forms the empirical foundation for our theoretical analysis in Section 5.
+4 MoLAE: Mixture of Latent Experts
+In this section, we introduce the Mixture of Latent Experts (MoLAE), a novel framework that maps experts into latent space to address several limitations of traditional MoE models.
+For clarity and focus, we exclude shared-experts from our analysis throughout this paper.
 4.1 Mixture of Latent Experts: Concept and Design
-To address these limitations, we propose the Mixture of Latent Experts (MoLE) framework, which fundamentally reimagines the structure of FFN layers in expert-based systems.
-Our approach is informed by a careful analysis of the expert computation in Equation equation 3.1, which comprises three distinct operations:
+To address these limitations, we propose MoLAE framework, which fundamentally redefines the structure of FFN layers in expert-based systems.
+Our approach is informed by a careful analysis of the expert computation in Equation (1), which comprises three distinct operations:
   1. 1.
@@
     Projection out: The intermediate output is mapped back from the low-dimensional space to the original high-dimensional space.
-A critical insight from this analysis is that the core functionality of experts primarily stems from the non-linear transformation in the second step.
-The projection operations in the first and third steps primarily serve to reduce computational complexity, potentially at the cost of limiting the expert’s specialized knowledge capacity.
-Drawing inspiration from both Multi-Head Latent Attention (MLA) (Liu et al., 2024), which introduces a “latent space” for KV caches in attention layers, and Grouped-Query Attention (GQA) (Ainslie et al., 2023), which leverages group-based processing, we propose the Mixture of Latent Experts (MoLE).
-This architecture fundamentally reconsiders how inputs are mapped to a lower-dimensional latent space where expert computation occurs more efficiently.
+A critical insight is that the core functionality of experts primarily stems from the non-linear transformation in the second step.
+The projection operations in the first and third steps primarily serve to reduce computational complexity, potentially at the cost of limiting the expert’s domain capacity.
+Drawing inspiration from both Multi-Head Latent Attention (MLA) (Liu et al., 2024), which introduces a “latent space” for KV caches in attention layers, and Grouped-Query Attention (GQA) (Ainslie et al., 2023), which leverages group-based processing, we propose the MoLAE, which operates on experts, i.e.
+specific FFN layers in standard MoE models.
+This architecture fundamentally reconsiders how inputs are projected into a lower-dimensional latent space, enabling more efficient computation within the experts.
 4.2 Formulation of Mixture of Latent Experts
@@
 Using the “up operator” $W_{\text{up}}^{i}$ of expert $i$ as a representative example, we propose a structured factorization where:
-$W_{\text{up}}^{i}=A_{\text{up}}^{i}B_{\text{up}}.$ (3.3)
+$W_{\text{up}}^{i}=A_{\text{up}}^{i}B_{\text{up}}.$ (3)
 In this formulation, $B_{\text{up}}\in\mathbb{R}^{m\times n}$ functions as a unified projection operator shared across experts, mapping inputs from the high-dimensional space $\mathbb{R}^{n}$ to a lower-dimensional latent space $\mathbb{R}^{m}$, where typically $m\ll n$.
 Conversely, $A_{\text{up}}^{i}\in\mathbb{R}^{m\times m}$ represents an expert-specific linear transformation within this latent space, encapsulating the specialized function of expert $i$.
-Following the terminology established in Multi-Head Latent Attention (MLA), we designate $B_{\text{up}}$ as the latent mapping for the “up operator”.
+Following the terminology established in MLA, we designate $B_{\text{up}}$ as the latent mapping for the “up operator”.
 This factorization principle extends systematically to the “gate operator” as well.
@@
 To optimize the trade-off between model expressivity and parameter efficiency, we introduce a structured grouping mechanism where each subset of $k$ experts shares the same latent mapping matrices $B_{\text{up}}$ and $B_{\text{down}}$.
-This design establishes a configurable spectrum of architectural possibilities: when $k=1$, each expert maintains its independent latent space, and MoLE becomes functionally equivalent to the standard MoE architecture.
+This design establishes a configurable spectrum of architectural possibilities: when $k=1$, each expert maintains its independent latent space, and MoLAE becomes functionally equivalent to the standard MoE architecture.
 Conversely, as $k$ increases, the model achieves progressively higher parameter efficiency at a measured trade-off with expert specialization.
 This parameterization allows for systematic exploration of the efficiency-performance frontier in mixture-of-experts architectures.
-3.4 Formal Definition of MoLE
 […]
-$E_{i}(x)=B_{\text{down}}^{\lfloor i/k\rfloor}A_{\text{down}}^{i}\left(A_{\text{up}}^{i}B_{\text{up}}^{\lfloor i/k\rfloor}(x)\odot\text{Act}\left(A_{\text{gate}}^{i}B_{\text{gate}}^{\lfloor i/k\rfloor}(x)\right)\right).$ (3.4)
-Consequently, the output of the FFN layer employing the MoLE architecture is computed as:
-$y=x+\sum_{i=1}^{N}g_{i}(x)E_{i}(x).$ (3.5)
+$E_{i}(x)=B_{\text{down}}^{\lfloor i/k\rfloor}A_{\text{down}}^{i}\left(A_{\text{up}}^{i}B_{\text{up}}^{\lfloor i/k\rfloor}(x)\odot\textsc{Act}\left(A_{\text{gate}}^{i}B_{\text{gate}}^{\lfloor i/k\rfloor}(x)\right)\right),$ (4)
+where $k$ is the group size of experts. Consequently, the output of the FFN layer employing the MoLAE architecture is computed as:
+$y=x+\sum_{i=1}^{N}g_{i}(x)E_{i}(x).$ (5)
 This formulation effectively disentangles expert-specific computations from the shared dimensionality reduction operations, enabling significant parameter efficiency while preserving model expressivity.
-The visual comparison between MoE and MoLE architectures is shown in Figure 1.
-3.5 Efficiency Benefits of MoLE
-The MoLE architecture offers significant efficiency advantages over standard fine-grained MoE models, spanning multiple computational dimensions from memory usage to communication overhead.
-To quantify these benefits, we provide a comparative analysis between MoE and MoLE for a single FFN layer, assuming identical configurations with hidden dimension $n$, MoE intermediate dimension $m$, and number of experts $N$.
+The visual comparison between MoE and MoLAE architectures is shown in Figure 1.
+4.3 Efficiency Benefits of MoLAE
+The MoLAE architecture offers significant efficiency advantages over standard MoE models, spanning multiple computational dimensions from memory usage to communication overhead.
+To quantify these benefits, we provide a comparative analysis between MoE and MoLAE for a single FFN layer, assuming identical configurations with hidden dimension $n$, MoE intermediate dimension $m$, and number of experts $N$.
+Table 2: Efficiency comparison between standard MoE and our proposed MoLAE architectures for a single FFN layer.
 Architecture Parameter Count FLOPs per Forward Pass
 Standard MoE $3Nmn$ $(3mn+2m)N$
-MoLE (Ours) $3Nm^{2}+3\lfloor\frac{N}{k}\rfloor mn$ $\left(3m^{2}+2m\right)N+3\lfloor\frac{N}{k}\rfloor mn$
-Table 1: Efficiency comparison between standard MoE and our proposed MoLE architectures for a single FFN layer.
-Parameter count represents the total number of trainable parameters in expert modules, while FLOPs measures the computational cost during a forward pass with top-1 routing.
-Here, $n$ represents the hidden dimension, $m$ the MoE intermediate dimension, $N$ the number of experts, and $k$ the number of latent experts per group.
-Parameter Efficiency As shown in Table 1, MoLE substantially reduces the parameter count compared to standard MoE, particularly when $m\ll n$, which is the typical case in modern LLMs.
+MoLAE (Ours) $3Nm^{2}+3\lfloor\frac{N}{k}\rfloor mn$ $\left(3m^{2}+2m\right)N+3\lfloor\frac{N}{k}\rfloor mn$
+Parameter Efficiency As shown in Table 2, MoLAE substantially reduces the parameter count compared to standard MoE, particularly when $m\ll n$, which is the typical case in modern LLMs.
 For instance, in DeepSeek-V3, $n=7168$ while $m=2048$.
 The parameter reduction stems from our latent parameterization, where expert-specific transformations operate in the lower-dimensional latent space ($m\times m$) rather than directly on the high-dimensional hidden space ($m\times n$).
-Computational Efficiency Beyond parameter savings, MoLE also reduces the computational cost measured in FLOPs.
+Computational Efficiency Beyond parameter savings, MoLAE also reduces the computational cost measured in FLOPs.
 The efficiency gain becomes particularly pronounced when the number of experts $N$ is large and $k$ is small (meaning fewer latent projection matrices are used).
 This computational advantage translates to faster inference and training times, especially on hardware where memory bandwidth is a bottleneck.
-Communication Overhead Reduction A critical but often overlooked benefit of MoLE is the reduction in all-to-all communication costs during distributed training and inference.
+Communication Overhead Reduction A critical but often overlooked benefit of MoLAE is the reduction in all-to-all communication costs during distributed training and inference.
 In standard MoE models, the full expert parameters ($3Nmn$ in total) must be synchronized across devices.
-In contrast, MoLE requires synchronization of significantly fewer parameters, reducing network bandwidth requirements and improving scalability for distributed deployments.
-Memory Access Patterns MoLE also offers improved cache efficiency during computation.
+In contrast, MoLAE requires synchronization of significantly fewer parameters, reducing network bandwidth requirements and improving scalability for distributed deployments.
+Memory Access Patterns MoLAE also offers improved cache efficiency during computation.
 The smaller matrices used in latent transformations ($A^{i}\in\mathbb{R}^{m\times m}$) exhibit better locality of reference compared to the larger matrices in standard MoE ($W^{i}\in\mathbb{R}^{m\times n}$), potentially leading to higher utilization of fast cache memory and reduced main memory bandwidth demands.
-4 Theoretical Analysis of MoE to MoLE Transformation
-In this section, we establish the theoretical foundation for transforming a standard Mixture of Experts (MoE) model into its corresponding Mixture of Latent Experts (MoLE) counterpart.
+5 Transformation from MoE to MoLAE
+In this section, we establish the theoretical foundation for transforming a standard MoE model into its corresponding MoLAE counterpart.
 We focus on the “up operator” as a representative example.
 For analytical clarity, we make two simplifying assumptions: (1) we omit the subscript of $W_{\text{up}}^{i}$ for ease of notation during this analysis, and (2) we consider the case where $k=N$, implying a single shared latent space operator for all experts.
-Keep the “down operator”.
-Note that for the “down operator” $W_{\text{down}}$, it maps the intermediate term from a low dimension into a high dimension, which captures the intrinsic properties of the given expert.
-Moreover, empirical experiments in Appendix C show that maintaining the structure of the “down operator” performs much better than not.
-Hence, we maintain the “down operator” in the transformed MoLE architecture and only change the “up operator” and “gate operator” from the MoE structure into the MoLE structure.
+Keep the “up operator”.
+Experiments on Qwen1.5-MoE-A2.7B indicate that transferring all operators during MoLAE transformation does not hurt performance.
+However, for Moonlight-16B-A3B, empirical results (Appendix B) show superior performance when preserving the “up operator” structure.
+Thus, for Moonlight, only its “down” and “gate” operators are converted to the MoLAE structure, retaining the “up operator”.
+Conversely, all operators are transferred for Qwen models.
 5.1 Transformation via Matrix Factorization
-For a given weight matrix $W^{i}$ associated with expert $i$, we aim to find corresponding matrices $A^{i}$ and $B$ such that $W^{i}x\approx A^{i}Bx$ for any input $x$.
-This transformation represents the post-training perspective where we seek to convert pre-trained MoE parameters into the MoLE architecture.
+For a given weight matrix $W^{i}$ associated with expert $i$, we aim to find corresponding matrices $A^{i}$ and $B$ such that $W^{i}X\approx A^{i}BX$ for the activation $X$.
+This transformation represents the post-training perspective where we seek to convert pre-trained MoE parameters into the MoLAE architecture.
 One direct approach is to determine matrices $A^{i}$ and $B$ such that $A^{i}B\approx W^{i}$ for all $i\in\{1,2,\ldots,N\}$. This leads naturally to the following optimization problem:
-$\min_{A^{i},B}\quad F(A^{1},\cdots,A^{N},B):=\frac{1}{2}\sum_{i=1}^{N}\|W^{i}-A^{i}B\|_{F}^{2},$ (4.1)
-where $\|\cdot\|_{F}$ denotes the Frobenius norm. Problem equation 4.1 admits the theoretical optimal solutions by using the SVD decomposition technique.
+$\min_{A^{i},B}\quad F(A^{1},\cdots,A^{N},B):=\frac{1}{2}\sum_{i=1}^{N}\|W^{i}-A^{i}B\|_{F}^{2},$ (6)
+where $\|\cdot\|_{F}$ denotes the Frobenius norm. Problem (6) admits the theoretical optimal solutions by using the SVD decomposition technique.
 Closed-Form Solution via Singular Value Decomposition.
-Obviously, problem equation 4.1 has infinitely many solutions since one can change one optimal solution $(A^{*},B^{*})$ into another by selecting a constant $\lambda$ and obtain another optimal solution $(\lambda A^{*},\frac{1}{\lambda}B^{*})$.
+Obviously, problem (6) has infinitely many solutions since one can change one optimal solution $(A^{*},B^{*})$ into another by selecting a constant $\lambda$ and obtain another optimal solution $(\lambda A^{*},\frac{1}{\lambda}B^{*})$.
 Hence, we only provide “one” optimal solution here.
-To derive a closed-form solution to problem equation 4.1, we consolidate the matrices into the following block structures:
-$W=\begin{pmatrix}W^{1}\\ W^{2}\\ \vdots\\ W^{N}\end{pmatrix},\quad A=\begin{pmatrix}A^{1}\\ A^{2}\\ \vdots\\ A^{N}\end{pmatrix}.$ (4.2)
-With this notation, problem equation 4.1 can be reformulated as:
-$\min_{A,B}\quad\frac{1}{2}\|W-AB\|_{F}^{2},$ (4.3)
+To derive a closed-form solution to problem (6), we consolidate the matrices into the following block structures:
+$W=\begin{pmatrix}W^{1}\\ W^{2}\\ \vdots\\ W^{N}\end{pmatrix},\quad A=\begin{pmatrix}A^{1}\\ A^{2}\\ \vdots\\ A^{N}\end{pmatrix}.$ (7)
+With this notation, problem (6) can be reformulated as:
+$\min_{A,B}\quad\frac{1}{2}\|W-AB\|_{F}^{2},$ (8)
 where $A\in\mathbb{R}^{mN\times m}$ and $B\in\mathbb{R}^{m\times n}$.
-While problem equation 4.3 has infinitely many solutions due to its underdetermined nature, the Eckart-Young-Mirsky theorem (Schmidt, 1989) provides an optimal solution with respect to the Frobenius norm.
+While problem (8) has infinitely many solutions due to its underdetermined nature, the Eckart-Young-Mirsky theorem (Schmidt, 1989) provides an optimal solution with respect to the Frobenius norm.
 Specifically, let $W=U\Sigma V^{\top}$ be the singular value decomposition (SVD) of $W$, where:
@@
 Let $\Sigma_{m}$ be the truncated version of $\Sigma$ that retains only the $m$ largest singular values: $\Sigma_{m}=\text{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{m},0,\ldots,0)\in\mathbb{R}^{mN\times n}$.
-According to the Eckart-Young-Mirsky theorem, an optimal solution to problem equation 4.3 is given by:
-$A^{*}=U_{m}\Sigma_{m}^{1/2},\quad B^{*}=\Sigma_{m}^{1/2}V_{m}^{\top},$ (4.4)
+According to the Eckart-Young-Mirsky theorem, an optimal solution to problem (8) is given by:
+$A^{*}=U_{m}\Sigma_{m}^{1/2},\quad B^{*}=\Sigma_{m}^{1/2}V_{m}^{\top},$ (9)
 where $U_{m}\in\mathbb{R}^{mN\times m}$ consists of the first $m$ columns of $U$, $\Sigma_{m}^{1/2}\in\mathbb{R}^{m\times m}$ is a diagonal matrix containing the square roots of the $m$ largest singular values, and $V_{m}\in\mathbb{R}^{n\times m}$ consists of the first $m$ columns of $V$.
 This factorization yields the minimum Frobenius norm error among all rank-$m$ approximations of $W$, with the approximation error given by:
-$\|W-A^{*}B^{*}\|_{F}^{2}=\sum_{i=m+1}^{\min(n,mN)}\sigma_{i}^{2}.$ (4.5)
-4.2 Minimizing Factorization Residuals
-In the previous subsection, we established methods for transforming MoE models into their MoLE counterparts from a theoretical perspective.
-Here, we address a critical aspect of this transformation: minimizing the residual error that inevitably arises when factorizing expert weights.
-We begin by establishing the precise conditions under which exact factorization is possible.
-Theorem 1.
-Given matrices $W^{i}\in\mathbb{R}^{m\times n}$ with $m\leq n$, there exist matrices $A^{i}\in\mathbb{R}^{m\times m}$ and a common matrix $B\in\mathbb{R}^{m\times n}$ such that $A^{i}B=W^{i}$ for all $i\in\{1,2,\ldots,N\}$, if and only if there exists an $(n-m)$-dimensional subspace $K\subseteq\mathbb{R}^{n}$ satisfying:
-$K\subseteq\bigcap_{i=1}^{N}\ker(W^{i}).$ (4.6)
-Proof.
-Necessity: Suppose there exist $B\in\mathbb{R}^{m\times n}$ and $A^{i}\in\mathbb{R}^{m\times m}$ such that $W^{i}=A^{i}B$ for all $i$.
-Since $m\leq n$ and we require exact factorization, $B$ must be row-full-rank (i.e., $\text{rank}(B)=m$).
-Consequently, its right nullspace $\ker(B)$ has dimension $n-m$.
-For any vector $x\in\ker(B)$, we have:
-$W^{i}x=A^{i}Bx=A^{i}\cdot 0=0,$ (4.7)
-which implies $\ker(B)\subseteq\ker(W^{i})$ for all $i$. Setting $K=\ker(B)$, we obtain an $(n-m)$-dimensional subspace contained in the intersection of all $\ker(W^{i})$.
-Sufficiency: Suppose there exists an $(n-m)$-dimensional subspace $K\subseteq\mathbb{R}^{n}$ such that $K\subseteq\ker(W^{i})$ for all $i$.
-We can construct $B\in\mathbb{R}^{m\times n}$ such that $\ker(B)=K$.
-Since $\dim(K)=n-m$, the matrix $B$ has rank $m$ by the rank-nullity theorem.
-For each $W^{i}$, the inclusion $K\subseteq\ker(W^{i})$ implies that any vector mapped to zero by $B$ is also mapped to zero by $W^{i}$.
-By the fundamental theorem of linear algebra, this means $\text{Row}(W^{i})\subseteq\text{Row}(B)$, where $\text{Row}(\cdot)$ denotes the row space.
-Therefore, there exists $A^{i}\in\mathbb{R}^{m\times m}$ such that $W^{i}=A^{i}B$ for each $i$.
-∎
-Theorem 1 provides a geometric interpretation of the factorization problem: exact factorization is possible only when the nullspaces of all expert matrices share a sufficiently large common subspace.
-In practical LLM implementations, however, this condition is rarely satisfied for FFN layers in MoE models, as our empirical analysis confirms.
-Given that exact factorization is generally unattainable, we now consider how to minimize the approximation error through strategic rank reduction.
-The rank-nullity theorem (Lang, 1987) states that for any linear mapping $W:X\rightarrow Y$:
-$\text{rank}(W)+\dim(\ker(W))=\dim(X).$ (4.8)
-In our context, $X$ represents the hidden space ($\mathbb{R}^{n}$) and $Y$ the MoE intermediate space ($\mathbb{R}^{m}$).
-Therefore, $\text{rank}(W^{i})+\dim(\ker(W^{i}))=n$ for all $i\in\{1,2,\ldots,N\}$.
-This relationship suggests a strategic approach: by reducing the rank of each $W^{i}$, we can increase the dimension of its nullspace.
-Specifically, if we constrain each $W^{i}$ to have rank $r<m$, then $\dim(\ker(W^{i}))=n-r>n-m$.
-This increases the probability of finding a substantial common subspace within the intersection $\bigcap_{i=1}^{N}\ker(W^{i})$, thereby improving the quality of our factorization.
-We implement this approach by computing low-rank approximations of each $W^{i}$ before attempting factorization.
-Importantly, our empirical experiments in Appendix A demonstrate that this rank reduction has minimal impact on model performance, suggesting that these FFN operators in MoE models inherently possess low-rank structure that can be exploited for more efficient parameterization.
-4.3 Transfer MoE to MoLE: A Unified Framework
-Based on our theoretical analyses in Sections 4.1 and 4.2, we now propose a unified, systematic framework for transforming Mixture of Experts (MoE) models into their more parameter-efficient Mixture of Latent Experts (MoLE) counterparts.
+$\|W-A^{*}B^{*}\|_{F}^{2}=\sum_{i=m+1}^{\min(n,mN)}\sigma_{i}^{2}.$ (10)
+To ensure the effectiveness of this transformation, we provide more details in Appendix A which discuss how to minimize the factorization residuals during the transformation.
+5.2 Transfer MoE to MoLAE: A Unified Framework
+Based on our theoretical analyses above, we propose a unified, systematic framework for transforming Mixture of Experts (MoE) models into their more parameter-efficient Mixture of Latent Experts (MoLAE) counterparts.
 Our framework consists of two principal steps, carefully designed to preserve model capabilities while enabling the latent parameterization.
@@
 For this step, we employ the SVD-based approach detailed in Section 5.1, which provides the optimal factorization with respect to the Frobenius norm.
-We formalize our approach in Algorithm 1, which provides a complete computational procedure for transforming MoE parameters into the MoLE architecture.
-Algorithm 1 Transformation of MoE to MoLE
+We formalize our approach in Algorithm 1, which provides a complete computational procedure for transforming MoE parameters into the MoLAE architecture.
+Algorithm 1 Transformation of MoE to MoLAE
 1:Expert weight matrices $\{W^{i}\}_{i=1}^{N}$, target rank $r$, latent dimension $m$
@@
 15:return $\{A^{i}\}_{i=1}^{N},B$
-This algorithm provides a computationally efficient procedure for transforming standard MoE layers into MoLE architecture.
+This algorithm provides a computationally efficient procedure for transforming standard MoE layers into MoLAE architecture.
 The rank reduction parameter $r$ and the latent dimension $m$ serve as hyperparameters that can be tuned to balance performance preservation against parameter efficiency.
@@
 Our framework is applicable to all linear operators within an MoE layer, including the up, down, and gate operators.
-By applying this transformation to each operator independently, we convert the entire MoE model to MoLE while maintaining model quality, as demonstrated in our experimental evaluations.
+By applying this transformation to each operator independently, we can convert the MoE model to MoLAE while maintaining overall performance.
 6 Experiments
-This section presents a comprehensive comparative analysis between Mixture of Latent Experts (MoLE) and Mixture of Experts (MoE) architectures in the context of pretraining GPT-2 (Radford et al., 2019) models and evaluating their subsequent performance on downstream tasks.
-5.1 Experimental Setup
-The MoE and MoLE models were configured with identical architectural parameters except for their Feed-Forward Network (FFN) layer structures, as detailed in Table 6.
+In this section, we evaluate the effectiveness of MoLAE on downstream tasks and the pre-training performance on GPT-2 (Radford et al., 2019).
+More experiments, like the importance of different “up, gate, down” operators and model configurations, are provided in Appendix B.
+Table 3: Comparisons of Qwen1.5-MoE-A2.7B and Moonlight-16B-A3B to their MoLAE architectures. To simplify the computations of efficiency, we only consider the MoE parts of the model.
+Models Benchmarks Params
+MMLU $\uparrow$ GSM8K $\uparrow$ CEval $\uparrow$ MNLI $\uparrow$ Wiki PPL $\downarrow$
+Qwen1.5-MoE 61.0 60.1 80.2 49.9 9.49 14.3B
+Qwen1.5-MoLAE
+($k$=10)
+60.0 59.4 74.1 40.6 9.74 10.5B
+[…]
+Moonlight-MoLAE
+($k$=8)
+60.2 72.3 62.8 38.7 10.6 12.3B
+6.1 Transformation from MoE to MoLAE: Downstream Tasks
+We first present our empirical analysis of transforming standard MoE architectures into their corresponding MoLAE counterparts.
+We specifically investigate two popular MoE models, including Qwen1.5-MoE-A2.7B model Team (2024) and Moonlight-16B-A3B (Liu et al., 2025), on diverse tasks, such as MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), CEval (Huang et al., 2023), MNLI (Wang et al., 2019) and Wikitext-2 (Merity et al., 2016).
+As shown in Table 3, MoLAE achieves performance comparable to that of the original MoE models, while exhibiting notable parameter efficiency and only slight performance degradations across the benchmarks.
+This result highlights the effectiveness of the transformation from MoE to MoLAE architectures.
+Furthermore, as indicated in Section 4, the group size $k$ in MoLAE architecture represents a critical hyperparameter that controls the trade-off between parameter efficiency and model expressivity.
+We evaluate several distinct configurations of $k\in\{1,10,20,30,60\}$ on the Qwen1.5-MoE model, which consists of 60 experts.
+Specifically, $k=1$ corresponds to the original MoE architecture; $k=10$ and $k=20$ represent balanced MoLAE configurations with multiple latent spaces; while $k=30$ and $k=60$ denote extreme cases with only one or two shared latent spaces, respectively.
+As shown in Table 2, with a moderate group size, such as $k=10$ for the Qwen1.5-MoE model, MoLAE largely preserves performance across various benchmarks.
+However, as the group size increases, the performance of the MoLAE model progressively deteriorates.
+For simpler tasks, such as CEval and MNLI, the performance decline is minimal, indicating that MoLAE retains most of the model’s capacity.
+In contrast, for more challenging tasks, such as GSM8K, there is a significant performance degradation.
+This dramatic decline empirically supports our theoretical analysis, which suggests that multiple latent spaces are essential to maintain the capacity of the original MoE model.
+Figure 2: Ablation study of group size $k$ on the Qwen1.5-MoE model.
+Figure 3: Comparison of training loss curves between MoE and MoLAE models on the English Wikipedia dataset.
+6.2 Pretraining of MoLAE
+To further validate the effectiveness of MoLAE, we construct the paired MoE and MoLAE models derived from the GPT-2 model (Radford et al., 2019), and then conduct the pretraining on the models.
+The MoE (151M) and MoLAE (94M) models are configured with identical architectural parameters except for FFN layer structures, as detailed in Table 5.
 Both architectures implement $N=32$ experts in their respective FFN layers.
-For the MoLE model, we establish $k=8$, indicating that each group of $8$ experts shares a common latent representation space.
-All models were trained on the Wikipedia English (Foundation, ) with sequences truncated to a maximum length of $512$ tokens.
-Training was conducted using the AdamW optimizer with consistent hyperparameters across all experimental conditions.
-The comprehensive specifications of model architectures and training configurations are presented in Table 6.
-5.2 Parameter Efficiency
-As quantified in Table 6, the introduction of shared latent spaces in MoLE architecture yields a substantial reduction in model parameter count compared to the standard MoE architecture.
-Specifically, MoLE achieves a 40% reduction in non-embedding parameters while maintaining comparable representational capacity.
+For MoLAE model, we establish $k=8$, indicating that each group of $8$ experts shares a common latent representation space.
+All models are trained on the Wikipedia English (Foundation, ) with maximum length of $512$.
+Models are updated using AdamW optimizer with consistent hyper-parameters across all runs.
+Parameter Efficiency
+As quantified in Table 5, the introduction of shared latent spaces in MoLAE architecture yields a substantial reduction in model parameter count compared to the standard MoE architecture.
+Specifically, MoLAE achieves a 40% reduction in non-embedding parameters while maintaining comparable model capacity.
 This parameter efficiency represents a significant advancement in model scalability without sacrificing performance.
-5.3 Training Dynamics
-Figure 2 illustrates the pretraining convergence characteristics of both MoE and MoLE models.
-The training loss trajectories reveal that MoLE maintains competitive optimization dynamics despite its significantly reduced parameter count.
-Although the MoE model converges to marginally lower loss values, this difference is negligible when considering the substantial parameter efficiency gained with MoLE.
-These results suggest that the shared latent representation in MoLE effectively preserves the essential modeling capacity while eliminating redundant parameterization inherent in traditional MoE architectures.
-Figure 2: Comparison of training loss curves between MoE and MoLE models on the English Wikipedia dataset.
-The convergence patterns demonstrate comparable performance, with MoLE exhibiting slightly higher loss values while utilizing substantially fewer parameters than the standard MoE architecture.
+Training Dynamics
+Figure 3 illustrates the pretraining convergence characteristics of both MoE and MoLAE models.
+The training loss trajectories reveal that MoLAE maintains competitive optimization dynamics despite its significantly reduced parameter count.
+Although the MoE model converges to marginally lower loss values, this difference is negligible when considering the substantial parameter efficiency gained with MoLAE.
+These results suggest that the shared latent representation in MoLAE effectively preserves the essential modeling capacity while eliminating redundant parameterization inherent in traditional MoE architectures.
+Using the well-trained MoE and MoLAE models, we evaluated their performance on the downstream task of Wikitext-2 perplexity.
+The results indicate that the performance gap between the two models is minimal, with PPL values of 79.5 and 81.5, respectively.
+This suggests that the MoLAE architecture serves as an effective pretraining base model, offering excellent efficiency while maintaining competitive performance.
 7 Conclusion
-In this paper, we presented Mixture of Latent Experts (MoLE), a novel parameter-efficient architecture that addresses the fundamental limitations of traditional MoE models.
-By factorizing expert weight matrices into shared projections and expert-specific transformations within a lower-dimensional latent space, MoLE significantly reduces parameter count and computational overhead without sacrificing model performance.
-Our comprehensive empirical evaluation demonstrates that MoLE not only preserves the capabilities of standard MoE models across diverse language tasks but also enables more economical scaling of large language models.
-The theoretical framework we developed for transforming pre-trained MoE models into the MoLE architecture provides a principled approach to expert parameterization, offering insights into the nature of redundancy in neural networks.
-As models continue to grow in scale and complexity, architectures like MoLE that prioritize parameter efficiency while maintaining performance will become increasingly valuable for practical deployment scenarios.
-Future work could explore extending these factorization techniques to other components of the transformer architecture and investigating dynamic adaptation of the latent space dimensionality based on input complexity.
-Appendix A Parameter Redundancy in Neural Networks
-This section provides empirical evidence for significant parameter redundancy within Feed-Forward Network (FFN) layers, thereby substantiating the theoretical framework presented in Section 4.
-We conduct our analysis using Qwen1.5-MoE-A2.7B [Team, 2024], a state-of-the-art Mixture-of-Experts architecture comprising $14.3$ billion total parameters while activating only $2.7$ billion parameters during inference.
-Each FFN layer in this architecture is characterized by an intermediate dimension $m=1408$, hidden dimension $n=2048$, and $N=60$ distinct expert modules.
-Figure 3: Singular value distributions of expert modules in the Qwen1.5-MoE-A2.7B model.
-The relatively uniform distributions indicate that weight matrices exhibit high-rank characteristics, with no significant concentration of information in a small subset of singular values.
-As illustrated in Figure 3, the singular value distributions of the expert modules demonstrate an absence of dominant singular values—a finding that contradicts conventional assumptions about the inherent low-rank structure of neural network weight matrices.
-This observation suggests that naive dimensional reduction techniques applying uniform low-rank approximations would potentially eliminate performance-critical parameters.
-Nevertheless, we demonstrate that despite the apparent full-rank nature of the weight matrices $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$, this characteristic is not a prerequisite for optimal model performance.
-Through systematic experimentation, we establish that these operators can be effectively replaced with carefully calibrated low-rank approximations while maintaining comparable performance metrics.
-For our analysis, we define a ratio-$r$ low-rank approximation $\tilde{W}$ of any matrix $W$ as a matrix whose rank equals $r\times\text{rank}(W)$, where $0<r\leq 1$.
-In accordance with the Eckart-Young-Mirsky theorem [Schmidt, 1989], these approximations are computed via Singular Value Decomposition (SVD), retaining only the largest $r\times\text{rank}(W)$ singular values and their corresponding singular vectors.
-To rigorously assess model capabilities under low-rank constraints, we evaluate performance on three benchmark tasks: MMLU [Hendrycks et al., 2021], GSM8K [Cobbe et al., 2021], and Wikitext-2 [Merity et al., 2016]—the same evaluation suite utilized in the original Qwen report.
-All experiments were conducted using the lm-evaluation-harness evaluation framework [Gao et al., 2024].
-Table 2 presents the comparative results.
-Low-rank ratio $r$ GSM8K (%) $\uparrow$ MMLU (%) $\uparrow$ Wikitext PPL $\downarrow$
-$1.0$ (Original) 60.1 61.0 9.49
-$0.8$ 61.2 60.3 9.65
-$0.6$ 60.1 59.4 9.85
-Table 2: Performance comparison of QwenMoE models with varying low-rank approximation ratios across multiple benchmarks.
-For the low-rank ratio $r=0.8$ case, the low-rank estimation model performs competitive with the original model on various tasks.
-Table 2 illustrates the relationship between rank reduction and model performance.
-The baseline case ($r=1.0$) represents the original, unmodified model with full-rank weight matrices.
-Notably, when reducing the rank of FFN operators by 20% ($r=0.8$), we observe no significant performance degradation.
-In fact, the reduced-rank model demonstrates superior performance on the GSM8K benchmark, achieving a 1.1 percentage point improvement over the original model, while maintaining comparable performance on MMLU accuracy and Wikitext-2 PPL.
-These empirical findings provide compelling evidence that, despite the mathematical full-rank property of FFN weight matrices, a substantial proportion of parameters contain redundant information that can be effectively approximated through lower-dimensional representations.
-This parameter redundancy phenomenon forms the empirical foundation for our theoretical analysis in Section 4.
-Appendix B Transformation from MoE to MoLE
-This section presents our empirical analysis of transforming standard MoE architectures into their corresponding MoLE counterparts.
-We specifically investigate the Qwen1.5-MoE-A2.7B model and systematically examine the impact of varying the latent space configuration parameter $k$ on model performance.
-As established in Section 3, the parameter $k$ in MoLE architecture represents a critical hyperparameter that controls the trade-off between parameter efficiency and model expressivity.
-We evaluate three distinct configurations of $k\in\{1,10,60\}$, where:
-  • $k=1$ corresponds to the original MoE architecture (baseline)
-  • $k=10$ represents a balanced MoLE configuration with multiple latent spaces
-  • $k=60$ represents the extreme case with a single shared latent space
-For consistent evaluation with Section A, we assess model performance on three established benchmarks: MMLU (reasoning), GSM8K (mathematical reasoning), and Wikitext-2 (language modeling perplexity).
-Table 3 presents the comparative performance across different MoLE configurations.
-Number of experts per latent ($k$) GSM8K (%) $\uparrow$ MMLU (%) $\uparrow$ Wikitext-2 PPL $\downarrow$
-$1$ (Original MoE) 60.1 61.0 9.49
-$10$ 59.4 60.0 9.74
-$60$ 14.3 40.1 25.97
-Table 3: Performance comparison of Qwen1.5-MoE-A2.7B with varying MoLE configurations.
-The balanced configuration ($k=10$) maintains near-parity with the original model while achieving significant parameter reduction.
-The single latent space configuration ($k=60$) demonstrates substantial performance degradation.
-The results demonstrate a clear relationship between latent space configuration and model performance.
-The balanced MoLE configuration ($k=10$) achieves comparable performance to the original MoE architecture, with only marginal degradation across all benchmarks: a 0.7 percentage point decrease on GSM8K, 1.0 percentage point decrease on MMLU, and 2.6% increase in perplexity on Wikitext-2.
-This configuration successfully reduces parameter count while preserving most of the model’s capabilities.
-In contrast, the extreme configuration ($k=60$) with a single shared latent space per FFN layer exhibits catastrophic performance degradation: a 45.8 percentage point decrease on GSM8K, 20.9 percentage point decrease on MMLU, and 173.7% increase in perplexity.
-This dramatic performance collapse empirically validates our theoretical analysis that multiple latent spaces are necessary to preserve the expressive capacity of the original MoE architecture.
-These findings provide compelling evidence that the MoLE architecture, when properly configured with an appropriate number of latent spaces, can achieve substantial parameter efficiency while maintaining competitive performance.
-Furthermore, they establish clear boundaries for the practical application of latent space factorization in expert-based neural architectures.
-Appendix C Critical Role of the “Down Operator”
-This section presents empirical evidence demonstrating that the “down operator” encapsulates more essential information than other components in Mixture of Experts (MoE) Large Language Models (LLMs).
-Through systematic experimentation, we establish the importance of preserving this operator’s structure for maintaining model performance.
-Consistent with our methodology in Appendix B, we utilize the Qwen1.5-MoE-A2.7B model with a fixed latent parameter $k=10$.
-To isolate the significance of the “down operator,” we implement two distinct transformation approaches:
+We introduce Mixture of Latent Experts (MoLAE), which overcomes limitations of traditional MoE models by factorizing expert weights into shared projections and expert-specific transformations in a lower-dimensional space.
+This approach reduces parameters and computation while maintaining performance across language tasks.
+Our theoretical framework for converting pre-trained MoE models to MoLAE provides insights into neural network redundancy.
+As models grow, such parameter-efficient architectures become increasingly valuable.
+Future work could extend these techniques to other transformer components and explore dynamic latent space adaptation.
+Appendix A Minimizing Factorization Residuals
+In the previous subsection 5, we established methods for transforming MoE models into their MoLAE counterparts from a theoretical perspective.
+Here, we address a critical aspect of this transformation: minimizing the residual error that inevitably arises when factorizing expert weights.
+We begin by establishing the precise conditions under which exact factorization is possible.
+Theorem 1
+Given matrices $W^{i}\in\mathbb{R}^{m\times n}$ with $m\leq n$, there exist matrices $A^{i}\in\mathbb{R}^{m\times m}$ and a common matrix $B\in\mathbb{R}^{m\times n}$ such that $A^{i}B=W^{i}$ for all $i\in\{1,2,\ldots,N\}$, if and only if there exists an $(n-m)$-dimensional subspace $K\subseteq\mathbb{R}^{n}$ satisfying:
+$K\subseteq\bigcap_{i=1}^{N}\ker(W^{i}).$ (11)
+Proof 1
+Necessity: Suppose there exist $B\in\mathbb{R}^{m\times n}$ and $A^{i}\in\mathbb{R}^{m\times m}$ such that $W^{i}=A^{i}B$ for all $i$.
+Since $m\leq n$ and we require exact factorization, $B$ must be row-full-rank (i.e., $\text{rank}(B)=m$).
+Consequently, its right nullspace $\ker(B)$ has dimension $n-m$.
+For any vector $x\in\ker(B)$, we have:
+$W^{i}x=A^{i}Bx=A^{i}\cdot 0=0,$ (12)
+which implies $\ker(B)\subseteq\ker(W^{i})$ for all $i$. Setting $K=\ker(B)$, we obtain an $(n-m)$-dimensional subspace contained in the intersection of all $\ker(W^{i})$.
+Sufficiency: Suppose there exists an $(n-m)$-dimensional subspace $K\subseteq\mathbb{R}^{n}$ such that $K\subseteq\ker(W^{i})$ for all $i$.
+We can construct $B\in\mathbb{R}^{m\times n}$ such that $\ker(B)=K$.
+Since $\dim(K)=n-m$, the matrix $B$ has rank $m$ by the rank-nullity theorem.
+For each $W^{i}$, the inclusion $K\subseteq\ker(W^{i})$ implies that any vector mapped to zero by $B$ is also mapped to zero by $W^{i}$.
+By the fundamental theorem of linear algebra, this means $\text{Row}(W^{i})\subseteq\text{Row}(B)$, where $\text{Row}(\cdot)$ denotes the row space.
+Therefore, there exists $A^{i}\in\mathbb{R}^{m\times m}$ such that $W^{i}=A^{i}B$ for each $i$.
+Theorem 1 provides a geometric interpretation of the factorization problem: exact factorization is possible only when the nullspaces of all expert matrices share a sufficiently large common subspace.
+In practical LLM implementations, however, this condition is rarely satisfied for FFN layers in MoE models, as our empirical analysis confirms.
+Given that exact factorization is generally unattainable, we now consider how to minimize the approximation error through strategic rank reduction.
+The rank-nullity theorem [Lang, 1987] states that for any linear mapping $W:X\rightarrow Y$:
+$\text{rank}(W)+\dim(\ker(W))=\dim(X).$ (13)
+In our context, $X$ represents the hidden space ($\mathbb{R}^{n}$) and $Y$ the MoE intermediate space ($\mathbb{R}^{m}$).
+Therefore, $\text{rank}(W^{i})+\dim(\ker(W^{i}))=n$ for all $i\in\{1,2,\ldots,N\}$.
+This relationship suggests a strategic approach: by reducing the rank of each $W^{i}$, we can increase the dimension of its nullspace.
+Specifically, if we constrain each $W^{i}$ to have rank $r<m$, then $\dim(\ker(W^{i}))=n-r>n-m$.
+This increases the probability of finding a substantial common subspace within the intersection $\bigcap_{i=1}^{N}\ker(W^{i})$, thereby improving the quality of our factorization.
+We implement this approach by computing low-rank approximations of each $W^{i}$ before attempting factorization.
+Importantly, our empirical experiments in Appendix 3.2 demonstrate that this rank reduction has minimal impact on model performance, suggesting that these FFN operators in MoE models inherently possess low-rank structure that can be exploited for more efficient parameterization.
+Appendix B Critical Role of the “Up Operator”
+While empirical experiments on Qwen1.5-MoE-A2.7B demonstrate that transferring all operators for MoE models can achieve superior performance, we observe that different operators contribute differentially to MoLAE transformation efficacy.
+This section presents empirical evidence establishing that the “up operator” encapsulates more essential information than other components in MoE models.
+Through systematic experimentation, we demonstrate the importance of preserving this operator’s structure for maintaining model performance.
+We examine a different MoE architecture, Moonlight-16B-A3B [Liu et al., 2025], in which the critical role of the “up operator” is more pronounced.
+Consistent with our methodology in Section 6.1, we utilize the Moonlight-16B-A3B model with a fixed latent parameter $k=8$, as it contains $64$ experts per layer.
+To isolate the significance of the “up operator,” we implement the following distinct transformation approaches:
   1. 1.
-    Partial transformation (“up+gate”) - Converts only the “up operator” and “gate operator” to their Mixture of Latent Experts (MoLE) equivalents while preserving the original “down operator.”
+    Partial transformation (“up+gate”) - Converts only the “up operator” and “gate operator” to theirMoLAE equivalents while preserving the original “down operator.”
   2. 2.
+    Partial transformation (“up+down”) - Converts only the “up operator” and “down operator” to theirMoLAE equivalents while preserving the original “gate operator.”
+  3. 3.
+    Partial transformation (“gate+down”) - Converts only the “gate operator” and “down operator” to theirMoLAE equivalents while preserving the original “up operator.”
+  4. 4.
     Complete transformation (“all”) - Transforms all three components (“up,” “gate,” and “down” operators) into their corresponding latent space representations.
-We evaluate these transformations across multiple downstream tasks, with results summarized in Table 4.
-Transformation Type ($k=10$) GSM8K (%) $\uparrow$ MMLU (%) $\uparrow$ Wikitext-2 PPL $\downarrow$
-Original MoE 60.1 61.0 9.49
-Partial (“up+gate”) 59.4 60.0 9.74
-Complete (“all”) 43.2 57.0 10.63
-Table 4: Performance comparison of Qwen1.5-MoE-A2.7B under different transformation configurations.
-Results demonstrate the critical importance of preserving the “down operator” structure for maintaining model performance.
-The results in Table 4 reveal a clear performance hierarchy.
-The original MoE model achieves optimal performance across all metrics.
-The partial transformation, which preserves the “down operator,” exhibits only marginal performance degradation (0.7% decrease on GSM8K, 1.0% on MMLU, and 0.25 increase in perplexity).
-In stark contrast, the complete transformation that modifies all operators, including the “down operator,” suffers substantial performance deterioration (16.9% decrease on GSM8K, 4.0% on MMLU, and 1.14 increase in perplexity).
-These findings provide compelling evidence that the “down operator” contains critical information that significantly impacts model performance.
+We evaluate these transformations on the MMLU and CEval [Huang et al., 2023] tasks, with results summarized in Table 4.
+Table 4: Performance comparison of Moonlight-16B-A3B under different transformation configurations.
+Results demonstrate the critical importance of preserving the “up operator” structure for maintaining model performance.
+Task ($k=8$) Original “up+gate” “up+down” “gate+down” “all”
+MMLU (%) $\uparrow$ 67.2 55.7 53.1 60.2 42.1
+CEval (%) $\uparrow$ 74.2 57.6 52.9 62.8 46.6
+In contrast to Qwen-MoE, the Moonlight-MoE architecture exhibits more substantial performance degradation under MoLAE transformation.
+As evidenced in Table 4, preserving the “up operator” yields optimal performance on downstream tasks.
+This phenomenon demonstrates the disproportionate importance of the “up operator” relative to the other components.
+These findings provide compelling evidence that the “up operator” encodes critical information that significantly influences model performance.
 When this operator is transformed into the latent space, substantial information loss occurs, resulting in markedly diminished capabilities across reasoning and knowledge-intensive tasks.
-This asymmetric importance among the operators suggests that architectural modifications to MoE models should prioritize preserving the structure of the “down operator” to maintain performance integrity.
-Appendix D Downstream Task Evaluation
-To rigorously assess the generalization capabilities of the proposed MoLE architecture, we conducted comprehensive evaluations across a diverse set of downstream tasks.
-The comparative performance metrics are presented in Table 5.
-Model Wikitext-2
-(PPL $\downarrow$)
-Standard MoE 75.86
-MoLE 81.57
-Table 5: Performance comparison between standard MoE and MoLE models across language modeling benchmark tasks. Bold values indicate superior performance.
-The empirical results in Table 5 demonstrate that despite utilizing approximately 40% fewer parameters than the standard MoE architecture, the proposed MoLE model achieves comparable performance across the PPL tasks.
-This parameter efficiency without substantial performance degradation highlights the effectiveness of our proposed architectural modification.
-Appendix E Training Arguments
-Hyperparameters MoE MoLE
+This asymmetric importance among operators suggests that architectural modifications to MoE models should prioritize preserving the structure of the “up operator” to maintain performance integrity.
+Appendix C Training Arguments
+See Table 5.
+Table 5: Model architecture and training hyperparameter configurations for MoE and MoLAE models.
+Hyperparameters MoE MoLAE
 FFN layers size $151$M $94$M
@@
 Learning rate schedule Cosine decay Cosine decay
-Table 6: Model architecture and training hyperparameter configurations for MoE and MoLE models.
+Appendix D Refined MoLAE Transformation
+In Section 5, we introduced the methodology to transform a standard MoE model into its corresponding MoLAE formulation.
+This section presents a refined approximation approach that incorporates activation information, resulting in enhanced precision.
+Given activation matrices $X^{i}$ for $i\in\{1,2,\ldots,N\}$, our objective is to determine low-rank factorization matrices $A^{i}$ and $B$ such that $W^{i}X^{i}\approx A^{i}BX^{i}$.
+While Section 5 assumed that the effect of activation matrices could be eliminated—simplifying the problem to Equation (6)—we now develop a more robust approximation that explicitly incorporates activation information.
+D.1 Problem Formulation
+We formulate the refined approximation as an optimization problem to find matrices $A^{i}$ and $B$ that minimize the sum of Frobenius norm differences between the original expert computations and their low-rank approximations:
+$\min_{A^{i},B}\quad F(A^{1},\ldots,A^{N},B):=\frac{1}{2}\sum_{i=1}^{N}\|W^{i}X^{i}-A^{i}BX^{i}\|_{F}^{2}$ (14)
+To facilitate the solution, we introduce block matrices $W$, $A$, and $X$ defined as:
+$W=\begin{pmatrix}W^{1}\\ W^{2}\\ \vdots\\ W^{N}\end{pmatrix},\quad A=\begin{pmatrix}A^{1}\\ A^{2}\\ \vdots\\ A^{N}\end{pmatrix},\quad X=\text{diag}(X^{1},X^{2},\ldots,X^{N})$ (15)
+This block representation transforms the problem in Equation (14) into an equivalent matrix factorization problem:
+$\min_{A,B}\quad\frac{1}{2}\|WX-ABX\|_{F}^{2}$ (16)
+This formulation enables us to derive an activation-aware low-rank approximation that more accurately preserves the input-output relationships of the original expert modules compared to the activation-agnostic approach described in Section 5.
+Problem (16) cannot be solved directly using the Eckart-Young-Mirsky theorem due to the presence of the activation matrix $X$.
+We therefore establish the following theorem that characterizes an optimal solution to problem (16).
+Theorem 2
+Let $X$ be the activation matrix and $W$ be the weight matrix.
+Assume that $X^{\top}X$ is positive definite with Cholesky decomposition $X^{\top}X=LL^{\top}$ where $L$ is invertible.
+Let $L^{\top}W=U\Sigma V^{\top}$ be the singular value decomposition of $L^{\top}W$ and $\Sigma_{m}$ be the truncated diagonal matrix containing the $m$ largest singular values, with corresponding truncated matrices $U_{m}$ and $V_{m}$.
+Then $A^{*}=(L^{\top})^{-1}U_{m}\Sigma_{m}^{1/2}$ and $B^{*}=\Sigma_{m}^{1/2}V_{m}^{\top}$ constitute an optimal solution to problem (16).
+Proof 2
+Since $X^{\top}X$ is positive definite, there exists a unique Cholesky decomposition $X^{\top}X=LL^{\top}$ where $L$ is invertible.
+Utilizing this decomposition, we can transform the original optimization problem as follows:
+$\|WX-ABX\|_{F}^{2}$ $=\text{Tr}[(WX-ABX)^{\top}(WX-ABX)]$
+$=\text{Tr}[(X^{\top}(W-AB)^{\top}(W-AB)X]$
+$=\text{Tr}[(W-AB)^{\top}XX^{\top}(W-AB)]$
+$=\text{Tr}[(W-AB)^{\top}LL^{\top}(W-AB)]$
+$=\text{Tr}[(L^{\top}(W-AB))^{\top}(L^{\top}(W-AB))]$
+$=\|L^{\top}W-L^{\top}AB\|_{F}^{2}$
+$=\|\tilde{W}-L^{\top}AB\|_{F}^{2},$
+where $\tilde{W}=L^{\top}W$.
+Given that $\tilde{W}=U\Sigma V^{\top}$ is the SVD of $\tilde{W}$, by the Eckart-Young-Mirsky theorem, the best rank-$m$ approximation of $\tilde{W}$ is $\tilde{W}_{m}=U_{m}\Sigma_{m}V_{m}^{\top}$, which can be factorized as $\tilde{W}_{m}=\tilde{A}\tilde{B}$ where $\tilde{A}=U_{m}\Sigma_{m}^{1/2}$ and $\tilde{B}=\Sigma_{m}^{1/2}V_{m}^{\top}$.
+We now demonstrate that $A^{*}=(L^{\top})^{-1}\tilde{A}$ and $B^{*}=\tilde{B}$ constitute an optimal solution to the original problem. First, for any matrices $A$ and $B$ with appropriate dimensions:
+$\min_{A,B}\|L^{\top}AB-\tilde{W}\|_{F}\leq\|L^{\top}A^{*}B^{*}-\tilde{W}\|_{F}=\|\tilde{A}\tilde{B}-\tilde{W}\|_{F}$ (17)
+Conversely, since $\tilde{A}\tilde{B}$ is the optimal rank-$m$ approximation of $\tilde{W}$:
+$\|\tilde{A}\tilde{B}-\tilde{W}\|_{F}=\min_{\text{rank}(T)\leq m}\|T-\tilde{W}\|_{F}\leq\|L^{\top}AB-\tilde{W}\|_{F},\quad\forall A,B$ (18)
+The final inequality holds because $\text{rank}(L^{\top}AB)\leq\text{rank}(AB)\leq m$ for any feasible solution $(A,B)$.
+Combining these inequalities:
+$\min_{A,B}\|L^{\top}AB-\tilde{W}\|_{F}=\|L^{\top}A^{*}B^{*}-\tilde{W}\|_{F}$ (19)
+Therefore, $(A^{*},B^{*})$ is an optimal solution to problem (16), which completes the proof.
+Based on Theorem 2, we propose the refined Algorithm 2 as follows.
+Algorithm 2 Refined MoLAE Transformation
+1:Weight matrices $W^{1},W^{2},\ldots,W^{N}$, activation matrices $X^{1},X^{2},\ldots,X^{N}$, and target rank $m$
+2:Low-rank factorization matrices $A^{1},A^{2},\ldots,A^{N}$ and $B$
+3:Construct block matrices $W=\begin{pmatrix}W^{1}\\ W^{2}\\ \vdots\\ W^{N}\end{pmatrix}$ and $X=\text{diag}(X^{1},X^{2},\ldots,X^{N})$
+4:Compute $X^{\top}X$
+5:if $X^{\top}X$ is singular then
+6:    Apply regularization: $X^{\top}X\leftarrow X^{\top}X+\lambda I$ for a small $\lambda>0$
+7:end if
+8:Compute the Cholesky decomposition: $X^{\top}X=LL^{\top}$
+9:Compute $\tilde{W}=L^{\top}W$
+10:Perform SVD on $\tilde{W}$: $\tilde{W}=U\Sigma V^{\top}$
+11:Extract the $m$ largest singular values and corresponding singular vectors:
+12:$\Sigma_{m}$, $U_{m}$, and $V_{m}$
+13:Compute $\tilde{A}=U_{m}\Sigma_{m}^{1/2}$
+14:Compute $B=\Sigma_{m}^{1/2}V_{m}^{\top}$
+15:Compute $A=(L^{\top})^{-1}\tilde{A}$
+16:Extract blocks of $A$ to obtain $A^{1},A^{2},\ldots,A^{N}$
+17:return $A^{1},A^{2},\ldots,A^{N},B$
+Appendix E Limitations
+We acknowledge certain limitations inherent to the present investigation.
+Constraints on available computational resources precluded evaluations of models at exceptionally large scales, notably DeepSeek-V3-671B and DeepSeek-R1-671B.
+Consequently, our empirical analysis was performed utilizing the Moonlight-16B-A3B model.
+This model employs an architectural design identical to that of DeepSeek-R1-671B, as elaborated upon in Appendix B.